### v2

# Target search optimization by threshold resetting

Abstract
We introduce a new class of first passage time optimization driven by threshold resetting, inspired by many natural processes where crossing a critical limit triggers failure, degradation or transition. In here, search agents are collectively reset when a threshold is reached, creating event-driven, system-coupled simultaneous resets that induce long-range interactions. We develop a unified framework to compute search times for these correlated stochastic processes, with ballistic- and diffusive searchers as key examples uncovering diverse optimization behaviors. A cost function, akin to breakdown penalties, reveals that optimal resetting can forestall larger losses. This formalism generalizes to broader stochastic systems with multiple degrees of freedom.
pacs:
Valid PACS appear here
Introduction. Resetting mediated target search has been a focal point of scientific interest in statistical physics and stochastic processes. A hallmark feature of resetting is that it can expedite the completion of complex processes by periodically reinstating a search process to a known or random state, thus uncovering new pathways, avoiding potential obstacles and in effect, enhancing the efficiency in locating desired targets evans_stochastic_2020; pal2024random. The advent of resetting based optimization techniques has truly invigorated much theoretical evans_diffusion_2011; kusmierz2014first; pal2016diffusion; reuveni_optimal_2016; pal_first_2017; chechkin2018random; belan2018restart; pal2019first; kumar2023universal; de2023resetting; pal2019landau; sar2023resetting; bonomo2021mitigating; de2022optimal; bhat2016stochastic; campos2015phase; pal_search_2020; biswas2024search; nagar2016diffusion, experimental tal2020experimental; besga2020optimal; faisant2021optimal; altshuler2024environmental; paramanick2024uncovering and computational jolakoski2023first; blumer2024combining; munoz2025learning research of first-passage phenomena.
The aim of this work is to introduce target search optimization under a threshold resetting (TR) mechanism. In contrast to the prototypical scenario where resetting is imposed by an external time clock, TR is an event driven phenomenon in which resetting occurs when the system operates under “safety covenant” or broadly, meets certain “threshold rules”. Thresholds play a crucial role in governing transitions across numerous systems, acting as critical points that trigger significant changes. In neuroscience, integrate-and-fire neurons accumulate stimuli until their membrane potential crosses a threshold, leading to an action potential before resetting to their initial state burkitt2006review; bachar2012stochastic; gerstein1964random; taillefumier2013phase. Financial strategies like stop-loss and take-profit rely on preset thresholds to limit losses or secure gains kaufman2013trading; shiryaev2007optimal; zhang2001stock; miller1966model. In software engineering, circuit breakers operate similar to threshold-based reset mechanisms, preventing an application to access an unresponsive server nygard2018release; surendro2021circuit; montesi2018decorator. In physics, fiber bundle models describe how materials distribute stress until a rupture threshold is met, after which the load is redistributed among the remaining fibers pradhan2010failure; hansen2015fiber. Simple stroboscopic threshold mechanisms have also been employed to obtain sustained temporal regularity in chaotic systems with applications to laser modeling sinha2001using; sinha1998dynamics; sinha1999computing; bhowmick2014targeting.
In the recent past, De Bruyne et al investigated spatial properties of a single diffusing particle under TR mechanism (referred as first-passage resetting there) disclosing dynamic nonstationarity of the process de2020optimization; de2021optimization. Yet, another unexplored aspect of the problem is the target search optimization – the study of which is at the heart of this letter. To put things in perspective, consider a first passage process that is conducted by a group of agents, setting off from a certain location and looking for a target until it is found redner2001; bray2013persistence; metzler_first-passage_2014; martinez2013optimizing; bhattacharya2014collective; falcon2019collective. The process is, however, restarted at some random time whenever one of the agents reaches a pre-assigned threshold. At that time all the searchers are reset simultaneously to the initial locations from which they renew their search. This collective resetting couples the degrees of freedom and makes the system strongly correlated biroli2023extreme; biroli2023critical; biroli2024exact. Such collective TR mechanism has natural justifications as earlier instances of experiments with fish school couzin2005effective; couzin2011uninformed; conradt2005consensus and swarming robots brambilla2013swarm; lerman2004review have demonstrated event-triggered collective transitions—where a local individual event (such as the detection of a fault or hazard by a single agent) prompts a coordinated adjustment in navigation strategy or “reset” across the population. Similarly, in individuals or groups of agents such as clonal raider ants (Ooceraea biroi), sensory thresholds shaped by factors like environment, and external stimuli influence decisions vital for survival and reproduction chase2025physics. Inspired by these connections, this letter aims to analyze the statistical signatures of the random completion time for this correlated multi-agent search process under threshold controlled resetting. To this end, we develop a unified framework for generic governing dynamics and show that TR can also be used as a potent mechanism to facilitate improved target search dynamics and efficiency. Finally, inspired from the control-management problem, we define a cost objective function that interpolates between the successful search time and the effective penalty owing to resetting events and examine whether this can be minimized.
General formalism. Consider a system of $N$ non-interacting searchers, each governed by the same stochastic dynamics which is generic, with position coordinates $\{x_{i}\}$, starting from $x_{0}$. The searchers are confined in an interval $[0,L]$ where origin is the target location. If any of the $N$ searchers reaches this target we mark the process complete. The TR mechanism is introduced by placing another boundary at $L$, which hereafter is referred to as the threshold (see Fig. 1). Whenever one of the searchers hits the threshold, all of them are collectively reset to the initial position $x_{0}$ from where the search is renewed – for a higher dimensional embedding interpretation, see  [^1] One natural way to interpret collective resetting is through the dynamics of a single ballistic searcher in $N$ dimensions, initialized at the point $(x_{0},x_{0},\ldots,x_{0})$. For simplicity, let us consider the case $N=2$. Imagine a square domain $(0,L)\times(0,L)$ with $0\leq x_{0}\leq L$. The target is located on two adjacent sides of the square: (1) $x_{2}=0$ with $0\leq x_{1}\leq L$, and (2) $x_{1}=0$ with $0\leq x_{2}\leq L$. The threshold, which triggers resetting, is defined by the other two sides: (3) $x_{1}=L$ with $0\leq x_{2}\leq L$, and (4) $x_{2}=L$ with $0\leq x_{1}\leq L$. A single particle begins at $(x_{0},x_{0})$ and moves with independently chosen velocities along the $x_{1}$ and $x_{2}$ directions. If it hits either of the threshold sides, it is reset to the initial point $(x_{0},x_{0})$ with new velocities — effectively resetting both coordinates simultaneously. If it hits one of the target sides, the process ends. This setup allows us to interpret the dynamics of two independent 1D walkers (in our model) as the two components of a single 2D walker undergoing simultaneous resetting. This interpretation generalizes naturally to higher dimensions: $N$ independent walkers correspond to the $N$ components of a single walker in $N$ dimensions, with the simultaneous resetting occurring when any coordinate reaches its respective threshold. This provides an intuitive and unified perspective on the simultaneous resetting mechanism explored in our work. See also Fig. S11. . This procedure repeats until the target at the origin is found, with the process considered complete as soon as any one searcher locates it. We show that despite the presence of strong correlations, the mean search/first-passage time (MFPT) can be exactly calculated and furthermore, a nontrivial optimization is discovered due to the threshold resetting and the emerging collective nature of the searchers. Note that the model assumes instantaneous information transfer among searchers: whenever one reaches the threshold, all reset simultaneously; likewise, when one finds the target, all are immediately informed and the search ends.
Figure 1: Schematic representation of the TR mechanism with $N$ searchers (here, $N=3$) starting from $x_{0}$ and undergoing generic stochastic dynamics. The process is considered complete when any of the searchers reaches the target located at the origin, and the corresponding first passage time is denoted by $\mathcal{T}_{N}^{\text{TR}}(L,x_{0})$. However, if any searcher reaches the resetting threshold at position $L$ before the target is found, all searchers are collectively returned to their initial positions, and the search process is restarted.
Figure 2: Panel (a): Variation of the scaled MFPT (8) with respect to $u$ and $N$. The plot shows a global minima in the MFPT at $u=1$ for any $N>1$. A cross section for $N=7$ is shown in panel (b). Panel (b) - central figure: Optimization of $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle$ as a function of $u$ (shown by the solid blue curve) – superimposed with individual components as given by the RHS of (11) namely $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$ (shown by orange dotted-dashed line), $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ (shown by green dashed line) and $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ (shown by red dotted line). Simulation results are shown by the cross markers demonstrating an excellent match. The accompanying subpanels showcase typical stochastic trajectories (with $x$ and $t$ standing for the position and time, respectively)
generated for different values of $u$ (marked by the circles) showing disparate behavior (global minima, local maxima & minima, divergence) as explained in the main text. In this analysis, we fix $x_{0}=0.5$ and vary $L$ within the range $[0.5,333.33]$ to explore the regime corresponding to $u\in(0.0015,1)$.
We start by introducing the survival probability under TR as $\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)$ which estimates the probability that none of the $N$-particles has found the target at the origin, starting from $x_{0}$, up to time $t$ although they may remain confined without hitting the resetting threshold or they can hit the threshold once or multiple times. Summing over all these possibilities, we can write the following renewal equation for the survival probability
$\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)=$
$q_{N}(L,x_{0},t)+\int_{0}^{t}dt^{\prime}j_{L,N}(L,x_{0},t^{\prime})\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t-t^{\prime}),$~~~~ (1)
where $q_{N}(L,x_{0},t)$ as the survival probability that none of the searchers has hit either of the boundaries at $x=0$ or $x=L$ up to time $t$ and $j_{L,N}(L,x_{0},t^{\prime})$ is the probability flux at the resetting threshold $L$ from any incoming searcher reaching at time $t^{\prime}<t$. Eq. (1) renders the following interpretation: Starting from $x_{0}$, the searchers can survive the target upto time $t$ in two ways. In one plausible set of events, the searchers survive without hitting any of the boundaries which occurs with the probability $q_{N}(L,x_{0},t)$ – this is the “no resetting” scenario and the first term on the RHS. In the second case, whenever any one of the $N$ searchers hits the resetting threshold for the very first time (say at $t^{\prime}$), all are instantaneously reset to the starting position $x_{0}$ renewing the next trial for the search. The contribution for any particle hitting the resetting threshold at $L$ is essentially the flux $j_{L,N}(L,x_{0},t^{\prime})$ followed by the survival probability $\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t-t^{\prime})$ for the remaining duration. Taking the Laplace transform of Eq. (1) with respect to $t$ and rearranging, we find
$\widetilde{\mathcal{Q}}_{N}^{\text{TR}}(L,x_{0},s)=\frac{\widetilde{q}_{N}(L,x_{0},s)}{1-\widetilde{j}_{L,N}(L,x_{0},s)},$ (2)
where we have defined $\widetilde{Z}(\cdot,s)=\int_{0}^{\infty}~dt~e^{-st}~Z(\cdot,t)$ as the Laplace transform of the function $Z(\cdot,t)$. Eq. (2) can be used to obtain all the moments of the completion times. For instance, the MFPT can be defined as
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle\equiv\int_{0}^{\infty}dt~\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)=\frac{\langle T_{N}(L,x_{0})\rangle}{\epsilon_{0}(L,x_{0},N)},$ (3)
where $\langle T_{N}(L,x_{0})\rangle=\int_{0}^{\infty}~dt~q_{N}(L,x_{0},t)$ is the mean fastest first passage time out of $N$ searchers to reach either the target or the resetting threshold and $\epsilon_{0}(L,x_{0},N)=1-\int_{0}^{\infty}~dt~j_{L,N}(L,x_{0},t)$ is the splitting probability that any of the $N$ searchers reaches the target first before hitting the threshold.
Since the searchers are non-interacting and identical in nature, we can write $q_{N}(L,x_{0},t)=\left[Q(L,x_{0},t)\right]^{N}$, where $Q(L,x_{0},t)$ is the survival probability of a single searcher, starting at $x_{0}\in[0,L]$ upto time $t$, with absorbing boundaries at both 0 and $L$ majumdar2020extreme. On the other hand, $j_{L,N}(L,x_{0},t)$ consists of current to the resetting threshold from a single searcher i.e. $j_{L,1}(L,x_{0},t)$ barring the rest N1N-1 of them which survive with probability $\left[Q(L,x_{0},t)\right]^{N-1}$ so that $j_{L,N}(L,x_{0},t)=Nj_{L,1}(L,x_{0},t)\left[Q(L,x_{0},t)\right]^{N-1}$ (the factor $N$ accounts for the fact that any of the $N$ searchers can contribute to $j_{L,1}(L,x_{0},t)$). Substituting these statistical metrics into Eq. (3), we arrive at
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\frac{\int_{0}^{\infty}dt~[Q(L,x_{0},t)]^{N}}{1-\int_{0}^{\infty}dt~Nj_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}},$ (4)
which is a general expression for the MFPT and a key result of this paper. Importantly, the derivation does not rely on specific dynamics of the searchers or target configurations, and moreover, the framework can be generalized to higher-dimensional search processes.
Ballistic search under TR mechanism. To illustrate the theoretical framework developed above, we examine the paradigmatic case of ballistic searchers with randomized velocity. Such models have been commonly used to describe navigation pattern for some foragers such as microzooplankton, jackals and marine predators reynolds2009optimising; sims2008scaling, to scrutinize optimal strategies for identifying randomly located target sites santos2004optimal; james2008optimizing and also in statistical physics villarroel2018continuous; biroli2024exact; martens2012probability; sprenger2023dynamics; bechinger2016active. Crucially, its analytical tractability offers valuable insights that enhance our understanding of TR-facilitated target search processes. Subsequently, we have extended our analysis to diffusive searchers.
We assume that each ballistic searcher moves with a random initial velocity drawn from a distribution $\phi(v)$. If the target is found a priori, the process therein ends. However, if the resetting threshold is encountered prior to the target, all the searchers are simultaneously reset to $x_{0}$ following which they draw new random velocities from the same distribution. As outlined above, we first derive the survival probability and the probability current for a single ballistic searcher (Sec. S1 in SI)
$Q(L,x_{0},t)$ $=\Phi\left(\frac{L-x_{0}}{t}\right)+\Phi\left(\frac{x_{0}}{t}\right),$ (5)
$~j_{L,1}(L,x_{0},t)$ $=\frac{L-x_{0}}{t^{2}}~\phi\left(\frac{L-x_{0}}{t}\right),$ (6)
where $\Phi(v)=\int_{0}^{v}dv^{\prime}\phi(v^{\prime})$; and substitute these into Eq. (4) to obtain the MFPT. For symmetric velocity distribution and $N=1$, this reads SI
$\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle=2L\int_{0}^{\infty}dv~\frac{\phi(v)}{v}.$ (7)
Remarkably, the MFPT does not depend on the starting point $x_{0}$ for any symmetric velocity distribution $\phi(v)$. Notably, the integral in Eq. (7) converges only when $\phi(v)\to 0$ as $v\to 0$ so that $\phi(v)/v$ becomes integrable around $v=0$. Simply put, if $\phi(v)$ does not vanish as $v\to 0$, there is a nonzero probability of observing realizations with searchers acquiring an almost vanishing velocity. These searchers will hit either the target or the resetting threshold in an exceedingly large timescale resulting in a divergent MFPT.
MFPT for multiple searchers ($N>1$): To continue the analysis for multiple searchers, we will assume that the velocities are chosen from a probability distribution given by $\phi(v)=\frac{1}{2v_{0}}e^{-|v|/v_{0}}$. While the MFPT for a single searcher is diverging, as evident from Eq. (7), the MFPT for the collective walkers e.g., $N\geq 2$ turns out to be a finite quantity. For the exponential kernel, the MFPT can be scaled as $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u\equiv\frac{x_{0}}{L})\rangle=\frac{v_{0}}{x_{0}}\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle={\cal F}\left(u,N\right)$, where $\mathcal{F}$ is the scaling function and depends only on $N$ and a dimensionless ratio $0\leq u\leq 1$. Computation of ${\cal F}(u,N)$ for $N>1$ is a challenging task that we undertake in SI, yielding
$\mathcal{F}(u,N)=\frac{\int_{0}^{\infty}dt\,\left[1-\frac{1}{2}\,e^{-u/t}-\frac{1}{2}\,e^{-(1-u)/t}\right]^{N}}{N\,u^{2}\,\int_{0}^{\infty}\frac{dt}{2t^{2}}\,e^{-u/t}\,\left[1-\frac{1}{2}\,e^{-u/t}-\frac{1}{2}\,e^{-(1-u)/t}\right]^{N-1}}\,.$ (8)
For $N=2$, a closed form expression can be obtained following a non-trivial computation (Sec. S3 in SI)
${\cal F}(u,2)=\frac{2}{u\,(3-2u)}\,\left[\ln 2-u\,\ln u-(1-u)\,\ln(1-u)\right],$ (9)
while for $N>2$, we can extract the following asymptotic forms which will be useful in later analysis
${\cal F}(u,N)\approx\begin{cases}&\hskip-8.5359pt\frac{a_{N}}{2^{N}-1}\,\frac{1}{u}\quad\quad\quad\quad\quad\quad\quad\quad\quad{\rm as}\quad u\to 0\,\vskip 5.69046pt\\ &\hskip-7.11317pta_{N}-N\,(1-u)\,\ln(1-u)\,\quad{\rm as}\quad u\to 1,\end{cases}$ (10)
where $a_{N}=\int_{0}^{\infty}\frac{dx}{x^{2}}\left(1-e^{-x}\right)^{N}$ is a constant. The scaling function ${\cal F}(u,N)$, evaluated as a function of $u$ for fixed $N$, diverges as $u\to 0$, and decreases monotonically with increasing $u$ when $N<4$. Interestingly, for $N\geq 4$, the function exhibits non-monotonic behavior: it first develops a local minimum, followed by a local maximum, and eventually approaches a global minimum as $u\to 1$, see Fig. 2(a) and a representative case for $N=7$ in Fig. 2(b) (solid curve).
Delving further, we observed that, in the absence of TR ($L\to\infty$ or equivalently $u=x_{0}/L\to 0$ limit), the FPT density decays as $\sim\frac{Nx_{0}}{2^{N}t^{2}}$, leading to a divergent MFPT for any $N$. In stark contrast, under optimal TR conditions at $u=1$, the distribution exhibits a faster decay, $f_{\mathcal{T}_{N}^{\text{TR}}}(x_{0},t)\sim t^{-N-1}$, ensuring a finite MFPT as $N\geq 2$ (see Sec. S8 for the derivation of the exact and universal FPT density that holds for arbitrary symmetric velocity distribution).
Optimization and limiting behavior of the MFPT.– To study the optimal behavior, we rewrite the MFPT by summing over the stochastic timescales
$\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle=\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle\times\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle+\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle,$~~ (11)
where $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle=\frac{1-\epsilon_{0}(u,N)}{\epsilon_{0}(u,N)}$ is the mean number of TR events upto the completion, and $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ is the mean conditional waiting time between consecutive TR events and $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ is the mean time for the first-passage to the target conditioned on no TR events. Numerical plots of these quantities are given in Fig. 2(b) (details in SI).
In the limit $u\to 1$ when the resetting threshold is at the closest proximity to $x_{0}$, the resetting events are frequent and we find $\langle{\mathcal{N}}_{\text{TR}}(u\to 1,N)\rangle=2^{N}-1$ SI. However, since all such resetting events consume almost no time with $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u\to 1)\rangle\to 0$, the cumulative contribution to the MFPT from TR events becomes almost negligible. Notably, the finite contribution to the MFPT in this case comes solely from those correlated events where all the ballistic searchers collectively sample negative velocities right after the resetting or at the inception to drift right toward the target, marking a successful completion. Thus, $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u\to 1)\rangle$ effectively contributes to the MFPT whose asymptotic form is given by Eq. (10). The typical trajectories in this limit (see the bottom-right panel in Fig. 2b) confirm this observation. Quite interestingly, this emergent global optimality at $u\to 1$ aligns with a widely used threshold control strategy in chaos theory sinha2001using; bhowmick2014targeting.
However, when the searchers start from a finite distance away from the resetting threshold ($u\lessapprox 1$), $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ become finite (recall that it is exactly zero for the $u\to 1$ limit) and $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$ remains large enough so that the first term in Eq. (11) dominates the MFPT leading to the local maxima observed in Fig. 2(b) (corresponding trajectories shown in the top-right figure for $u=0.9$). Further tuning initial configuration of the searchers in the direction of the target (i.e. reducing $u$ further) ensures $\epsilon_{0}>1/2$ (see Fig. (S2) in SI) and subsequently, we see a gradual decrease in the MFPT upto a local minimum. Stochastic realizations at the bottom-left of Fig. 2(b) depict this scenario.
Moreover, in the limit $u\to 0$ where the threshold is far away from the starting position, each searcher can acquire a positive or negative velocity with probability $1/2$ for the symmetric velocity distribution and thus, there is a finite probability $\frac{1}{2^{N}}$ that all the $N$-searchers possess a positive velocity that is away from the target. The probability of selecting a unfavorable positive velocity, albeit vanishingly small, remains finite and yield in MFPT with a diverging form $\sim 1/u$ given by Eq. (10) (Sec. S5 in SI). Contributing realizations (for instance, shown in the top-left panel in Fig. 2b for $u=0.0015$) corroborate this observation.
Finally, examining the MFPT with respect to $N$, we find that depending on $u$ and the interaction mediated by $N$, this variation can be either monotonic or non-monotonic. Nevertheless, for all $N$, the MFPT attains its minimum at $u=1$ (see Fig. S2), further supporting our claim in the ballistic search scenario.
Figure 3: Behavior of the cost function with respect to $u=x_{0}/L$ for $N=3$ and $\beta=1$. The cost attains a minimum at an optimal $u^{*}$ – the variation of which as a function of NN is shown in the inset panel. The squares represent data from simulations.
Cost of TR mechanism. A central quest in the resetting discipline is to harness the cost, be dynamic de2020optimization; sunil2023cost; sunil2024minimizing or thermodynamic tal2020experimental; tal2025smart; olsen2024thermodynamic; gupta2020work; pal2023thermodynamic, of the overall process. Imitating the penalty during a power grid breakdown ajjarapu1994optimal; feng2000comprehensive, we posit, for each resetting event, a cost that is proportional to the MFPT as well as the total number of TR events. Henceforth, the goal is to determine the optimal operation of this system, for which the cost function
$C_{N}(u)=\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle+\beta N\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle,$ (12)
is minimal. Here, $\beta$ is a constant that can be thought of cost per resetting event for a single particle and thus, we multiply by $N$ to enumerate the total cost for the “reset operation”. For the ballistic searchers, Fig. 3 shows that $C_{N}(u)$ attains a minimum for an optimal $u=u^{*}$ for a fixed $N$. Note that unlike the MFPT, the cost function is not globally minimized at $u\to 1$ since in this limit the resetting frequency is significantly high and the second term in Eq. (12) dominates. On the other hand, if the resetting threshold is far away from the starting position, the MFPT gradually increases, contributing to diverging cost. In the inset of Fig. 3, we show the variation of $u^{*}$ with respect to $N$, further noting that large number of searchers can systematically reduce the optimal distance $u^{*}\to 0$. Interestingly, the overall cost function remains optimized despite the increase in individual cost through $\beta$, accompanied by a gradual rise in $u^{*}$ SI. Furthermore, $u^{*}$ decreases with increasing $\beta$ for a fixed $N$, thus reflecting the high cost of frequent resets SI. Evidently, this collective cost optimization is different than that of the MFPT, indicating that a judicious choice of optimal $u$ can leverage to a decision based biasing between the MFPT and the cost function.
Conclusions. Motivated by threshold-driven phenomena that refer to processes or events that occur only when a system surpasses a critical threshold, we study the first passage properties of an arbitrary search dynamics under threshold resetting mechanism. We develop a general theoretical framework and study the optimization properties of the MFPT which is a useful metric in learning the efficiency of a target search process. The optimization reported here also resonates with the performance maximization by the agents (robots or foraging animals) that collaborate and achieve tasks (for instance finding the desired target between many competitive ones by a fish school in couzin2011uninformed) in an optimal way.
In a classical foraging model where the searchers move ballistically, we discover a non-trivial optimization that suggests that persistent resetting to the threshold can render the MFPT globally minimum – an effect reminiscent to the threshold mechanism in nonlinear chaotic systems sinha2001using; bhowmick2014targeting. Extending the analysis to $N$ diffusive searchers also reveals rich optimization behaviors of the MFPT SI; diffusive_TR. By defining a cost function for such systems, we further demonstrate that the collective TR mechanism can achieve optimal cost.
To extend the formalism to capture more realistic scenarios, we can incorporate an overhead time for information transfer. This can be done in two ways: First, we can add a random (or deterministic) information passing time $\mathcal{T}^{O}$ whenever a searcher hits the threshold, after which all of them are reset to the initial position. In this case, one can show that the overall MFPT of the system would increase (however, this computation excludes the time required to transfer information to the others once a searcher successfully locates the target) diffusive_TR. A more involved extension would be to consider space–time coupled return protocols, where after a TR event searchers return to the initial configuration via finite-time spatial dynamics pal_search_2020; tal2020experimental; biswas2024search.
Finally, it is worth noting that the formalism can also be applied to extract higher order moments of the search time in higher dimensions and to the dynamics of non-coordinate variables like a fluctuating potential as in the case of firing neurons where the membrane potential of an individual synapse can reset after reaching a critical value burkitt2006review; taillefumier2013phase. Ongoing works also include characterizing nonequilibrium stationary states for such long-range systems under TR diffusive_TR; BMS2025. Given its simple implementation yet achieving target functionality makes the TR mechanism a promising avenue in the target search problem which can have far reaching implications beyond statistical physics to foraging, control theory, finance, structural failure-repair analysis and in computing systems.
Acknowledgments. SNM thanks M. Biroli, S. Redner and G. Schehr for discussions on related models. AP sincerely thanks D. Ghosh for pointing towards some relevant works. The numerical calculations reported in this work were carried out on the Kamet cluster, which is maintained and supported by the Institute of Mathematical Science’s High-Performance Computing Center. AB and AP gratefully acknowledge research support from the Department of Atomic Energy, Government of India. SNM acknowledges support from ANR Grant No. ANR- 23- CE30-0020-01 EDIPS. SNM and AP thank the Higgs Center for Theoretical Physics, Edinburgh, for hospitality during the workshop “New Vistas in Stochastic Resetting” and Korea Institute for advanced Study(KIAS), Seoul, for hospitality during the conference “Nonequilibrium Statistical Physics of Complex Systems” where several discussions related to the project took place. SNM and AP also acknowledge the research support from the International Research Project (IRP) titled “Classical and quantum dynamics in out of equilibrium systems” by CNRS, France.
End Matter
Diffusive searchers under TR. Let us consider the case of $N$ diffusive searchers with the same set-up as for the ballistic walkers. To proceed, we will need the single searcher diffusive propagator $G(x,t|x_{0})$ in the presence of two absorbing boundaries at $x=0$ and $x=L$. This is well known in literature and is given by redner2001
$G(x,t|x_{0})=\frac{2}{L}\sum_{n=0}^{\infty}\sin\left(\frac{\pi nx}{L}\right)\sin\left(\frac{\pi nx_{0}}{L}\right)e^{-\frac{n^{2}\pi^{2}Dt}{L^{2}}},$ (13)
where $D$ is the diffusion constant. Using this, we can calculate the survival probability and the currents through the threshold and the target in terms of two quantities $u$ and $\tau_{d}=L^{2}/D$ so that SI
$Q(u,t)=\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (14)
$j_{0,1}(u,t)=\frac{2\pi}{\tau_{d}}\sum_{n=1}^{\infty}n\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (15)
$j_{L,1}(u,t)=-\frac{2\pi}{\tau_{d}}\sum_{n=1}^{\infty}n(-1)^{n}\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (16)
Figure 4: Diffusive searchers under TR: Mean first-passage time (MFPT) as a function of the threshold parameter $u$ for different values of NN as in Eq. (17). The horizontal dashed lines (for $N\geq 3$) represent the MFPT in the absence of threshold resetting, corresponding to the $u\to 0$ limit (note that the MFPTs for $N=1,2$ in this limit are diverging). The fact that the MFPT curves fall below these baselines highlights the enhanced search efficiency achieved through the TR mechanism, also in the case of diffusive searchers. The markers represent the simulation results.
The MFPT can be written in a dimensionless form as $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle=\langle\mathcal{T}_{N}^{\text{TR}}(x_{0},L)\rangle/\tau_{d}=\mathcal{F}(u,N)$. Skipping details from SI, we obtain an exact expression of the scaling function $\mathcal{F}(u,N)$ representing the MFPT which is given by
$\mathcal{F}(u,N)$ $=\frac{1}{u^{2}}\left(\frac{\int_{0}^{\infty}dt~\left[\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-n^{2}\pi^{2}t}\right]^{N}}{N\int_{0}^{\infty}dt~\left[2\pi\sum_{n=1}^{\infty}n\sin\left(n\pi u\right)e^{-n^{2}\pi^{2}t}\right]\left[\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-n^{2}\pi^{2}t}\right]^{N-1}}\right),$ (17)
which we plot in Fig. 4 as a function of $u$ for different $N$. Summarizing the results from the plot and subsequent analysis SI, we note that for a single diffusive searcher, the MFPT decreases as $u^{-1}$. It is the lowest when $L\to x_{0}$ ($u\to 1$) where $\langle\overline{\mathcal{T}}_{1}^{\text{TR}}(u=1)\rangle\rangle=1/2$. However, for $N\geq 2$, the MFPT exhibits distinct optimization features. It should be noted that for $N=1$ and $N=2$, the TR mechanism always reduces the search time. It is observed from Fig. 4 that TR can facilitate the search process while compared to the same in the absence of TR as for intermediate values of $u$, the MFPT curves go below their respective dashed lines—each representing the reset-free case for different NN—clearly indicating that threshold resetting enhances search efficiency. Although it is known that the MFPT for diffusive searchers is finite only for $N\geq 3$ (see lindenberg1980lattice; krapivsky2010maximum; SI), the introduction of the TR mechanism enables a further reduction by optimally constraining the searchers’ trajectories, leading to an optimal resetting threshold in the range $0<u<1$ for all $N\geq 2$. This optimization observed in the diffusive case, alongside similar results for ballistic searchers, demonstrates that TR is a robust, event-driven mechanism capable of significantly accelerating target search in complex collective systems.