### v3

# Valid PACS appear here

Abstract
We introduce a new class of first passage time optimization driven by threshold resetting, inspired by many natural processes where crossing a critical limit triggers failure, degradation or transition. In here, search agents are collectively reset when a threshold is reached, creating event-driven, system-coupled simultaneous resets that induce long-range interactions. We develop a unified framework to compute search times for these correlated stochastic processes, with ballistic- and diffusive searchers as key examples uncovering diverse optimization behaviors. A cost function, akin to breakdown penalties, reveals that optimal resetting can forestall larger losses. This formalism generalizes to broader stochastic systems with multiple degrees of freedom.
pacs:
Valid PACS appear here
Introduction. Resetting mediated target search has been a focal point of scientific interest in statistical physics and stochastic processes. A hallmark feature of resetting is that it can expedite the completion of complex processes by periodically reinstating a search process to a known or random state, thus uncovering new pathways, avoiding potential obstacles and in effect, enhancing the efficiency in locating desired targets Evans et al. (2020); Pal et al. (2024). The advent of resetting based optimization techniques has truly invigorated much theoretical Evans and Majumdar (2011); Kusmierz et al. (2014); Pal et al. (2016); Reuveni (2016); Pal and Reuveni (2017); Chechkin and Sokolov (2018); Belan (2018); Pal et al. (2019); Kumar and Pal (2023); De Bruyne and Mori (2023); Pal and Prasad (2019); Sar et al. (2023); Bonomo et al. (2022); De Bruyne et al. (2022); Bhat et al. (2016); Campos and Méndez (2015); Pal et al. (2020); Biswas et al. (2024); Nagar and Gupta (2016), experimental Tal-Friedman et al. (2020); Besga et al. (2020); Faisant et al. (2021); Altshuler et al. (2024); Paramanick et al. (2024) and computational Jolakoski et al. (2023); Blumer et al. (2024); Muñoz-Gil et al. (2025) research of first-passage phenomena.
The aim of this work is to introduce target search optimization under a threshold resetting (TR) mechanism. In contrast to the prototypical scenario where resetting is imposed by an external time clock, TR is an event driven phenomenon in which resetting occurs when the system operates under “safety covenant” or broadly, meets certain “threshold rules”. Thresholds play a crucial role in governing transitions across numerous systems, acting as critical points that trigger significant changes. In neuroscience, integrate-and-fire neurons accumulate stimuli until their membrane potential crosses a threshold, leading to an action potential before resetting to their initial state Burkitt (2006); Bachar et al. (2012); Gerstein and Mandelbrot (1964); Taillefumier and Magnasco (2013). Financial strategies like stop-loss and take-profit rely on preset thresholds to limit losses or secure gains Kaufman (2013); Shiryaev (2007); Zhang (2001); Miller and Orr (1966). In software engineering, circuit breakers operate similar to threshold-based reset mechanisms, preventing an application to access an unresponsive server Nygard (2018); Surendro et al. (2021); Montesi and Weber (2018). In physics, fiber bundle models describe how materials distribute stress until a rupture threshold is met, after which the load is redistributed among the remaining fibers Pradhan et al. (2010); Hansen et al. (2015). Simple stroboscopic threshold mechanisms have also been employed to obtain sustained temporal regularity in chaotic systems with applications to laser modeling Sinha (2001); Sinha and Ditto (1998, 1999); Bhowmick et al. (2014).
In the recent past, De Bruyne et al investigated spatial properties of a single diffusing particle under TR mechanism (referred as first-passage resetting there) disclosing dynamic nonstationarity of the process De Bruyne et al. (2020); De Bruyne and Redner (2021). Yet, another unexplored aspect of the problem is the target search optimization – the study of which is at the heart of this letter. To put things in perspective, consider a first passage process that is conducted by a group of agents, setting off from a certain location and looking for a target until it is found Redner (2001); Bray et al. (2013); Metzler et al. (2014); Martínez-García et al. (2013); Bhattacharya and Vicsek (2014); Falcón-Cortés et al. (2019). The process is, however, restarted at some random time whenever one of the agents reaches a pre-assigned threshold. At that time all the searchers are reset simultaneously to the initial locations from which they renew their search. This collective resetting couples the degrees of freedom and makes the system strongly correlated Biroli et al. (2023a, b, 2024). Such collective TR mechanism has natural justifications as earlier instances of experiments with fish school Couzin et al. (2005, 2011); Conradt and Roper (2005) and swarming robots Brambilla et al. (2013); Lerman et al. (2004) have demonstrated event-triggered collective transitions—where a local individual event (such as the detection of a fault or hazard by a single agent) prompts a coordinated adjustment in navigation strategy or “reset” across the population. Similarly, in individuals or groups of agents such as clonal raider ants (Ooceraea biroi), sensory thresholds shaped by factors like environment, and external stimuli influence decisions vital for survival and reproduction Chase and Peleg (2025). Inspired by these connections, this letter aims to analyze the statistical signatures of the random completion time for this correlated multi-agent search process under threshold controlled resetting. To this end, we develop a unified framework for generic governing dynamics and show that TR can also be used as a potent mechanism to facilitate improved target search dynamics and efficiency. Finally, inspired from the control-management problem, we define a cost objective function that interpolates between the successful search time and the effective penalty owing to resetting events and examine whether this can be minimized.
General formalism. Consider a system of $N$ non-interacting searchers, each governed by the same stochastic dynamics which is generic, with position coordinates $\{x_{i}\}$, starting from $x_{0}$. The searchers are confined in an interval $[0,L]$ where origin is the target location. If any of the $N$ searchers reaches this target we mark the process complete. The TR mechanism is introduced by placing another boundary at $L$, which hereafter is referred to as the threshold (see Fig. 1). Whenever one of the searchers hits the threshold, all of them are collectively reset to the initial position $x_{0}$ from where the search is renewed – for a higher dimensional embedding interpretation, see  [^1] One natural way to interpret collective resetting is through the dynamics of a single ballistic searcher in $N$ dimensions, initialized at the point $(x_{0},x_{0},\ldots,x_{0})$. For simplicity, let us consider the case $N=2$. Imagine a square domain $(0,L)\times(0,L)$ with $0\leq x_{0}\leq L$. The target is located on two adjacent sides of the square: (1) $x_{2}=0$ with $0\leq x_{1}\leq L$, and (2) $x_{1}=0$ with $0\leq x_{2}\leq L$. The threshold, which triggers resetting, is defined by the other two sides: (3) $x_{1}=L$ with $0\leq x_{2}\leq L$, and (4) $x_{2}=L$ with $0\leq x_{1}\leq L$. A single particle begins at $(x_{0},x_{0})$ and moves with independently chosen velocities along the $x_{1}$ and $x_{2}$ directions. If it hits either of the threshold sides, it is reset to the initial point $(x_{0},x_{0})$ with new velocities — effectively resetting both coordinates simultaneously. If it hits one of the target sides, the process ends. This setup allows us to interpret the dynamics of two independent 1D walkers (in our model) as the two components of a single 2D walker undergoing simultaneous resetting. This interpretation generalizes naturally to higher dimensions: $N$ independent walkers correspond to the $N$ components of a single walker in $N$ dimensions, with the simultaneous resetting occurring when any coordinate reaches its respective threshold. This provides an intuitive and unified perspective on the simultaneous resetting mechanism explored in our work. See also Fig. S11. . This procedure repeats until the target at the origin is found, with the process considered complete as soon as any one searcher locates it. We show that despite the presence of strong correlations, the mean search/first-passage time (MFPT) can be exactly calculated and furthermore, a nontrivial optimization is discovered due to the threshold resetting and the emerging collective nature of the searchers. Note that the model assumes instantaneous information transfer among searchers: whenever one reaches the threshold, all reset simultaneously; likewise, when one finds the target, all are immediately informed and the search ends.
Figure 1: Schematic representation of the TR mechanism with $N$ searchers (here, $N=3$) starting from $x_{0}$ and undergoing generic stochastic dynamics. The process is considered complete when any of the searchers reaches the target located at the origin, and the corresponding first passage time is denoted by $\mathcal{T}_{N}^{\text{TR}}(L,x_{0})$. However, if any searcher reaches the resetting threshold at position $L$ before the target is found, all searchers are collectively returned to their initial positions, and the search process is restarted.
Figure 2: Panel (a): Variation of the scaled MFPT (8) with respect to $u$ and $N$. The plot shows a global minima in the MFPT at $u=1$ for any $N>1$. A cross section for $N=7$ is shown in panel (b). Panel (b) - central figure: Optimization of $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle$ as a function of $u$ (shown by the solid blue curve) – superimposed with individual components as given by the RHS of (11) namely $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$ (shown by orange dotted-dashed line), $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ (shown by green dashed line) and $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ (shown by red dotted line). Simulation results are shown by the cross markers demonstrating an excellent match. The accompanying subpanels showcase typical stochastic trajectories (with $x$ and $t$ standing for the position and time, respectively)
generated for different values of $u$ (marked by the circles) showing disparate behavior (global minima, local maxima & minima, divergence) as explained in the main text. In this analysis, we fix $x_{0}=0.5$ and vary $L$ within the range $[0.5,333.33]$ to explore the regime corresponding to $u\in(0.0015,1)$.
We start by introducing the survival probability under TR as $\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)$ which estimates the probability that none of the $N$-particles has found the target at the origin, starting from $x_{0}$, up to time $t$ although they may remain confined without hitting the resetting threshold or they can hit the threshold once or multiple times. Summing over all these possibilities, we can write the following renewal equation for the survival probability
$\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)=$
$q_{N}(L,x_{0},t)+\int_{0}^{t}dt^{\prime}j_{L,N}(L,x_{0},t^{\prime})\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t-t^{\prime}),$~~~~ (1)
where $q_{N}(L,x_{0},t)$ as the survival probability that none of the searchers has hit either of the boundaries at $x=0$ or $x=L$ up to time $t$ and $j_{L,N}(L,x_{0},t^{\prime})$ is the probability flux at the resetting threshold $L$ from any incoming searcher reaching at time $t^{\prime}<t$. Eq. (1) renders the following interpretation: Starting from $x_{0}$, the searchers can survive the target upto time $t$ in two ways. In one plausible set of events, the searchers survive without hitting any of the boundaries which occurs with the probability $q_{N}(L,x_{0},t)$ – this is the “no resetting” scenario and the first term on the RHS. In the second case, whenever any one of the $N$ searchers hits the resetting threshold for the very first time (say at $t^{\prime}$), all are instantaneously reset to the starting position $x_{0}$ renewing the next trial for the search. The contribution for any particle hitting the resetting threshold at $L$ is essentially the flux $j_{L,N}(L,x_{0},t^{\prime})$ followed by the survival probability $\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t-t^{\prime})$ for the remaining duration. Taking the Laplace transform of Eq. (1) with respect to $t$ and rearranging, we find
$\widetilde{\mathcal{Q}}_{N}^{\text{TR}}(L,x_{0},s)=\frac{\widetilde{q}_{N}(L,x_{0},s)}{1-\widetilde{j}_{L,N}(L,x_{0},s)},$ (2)
where we have defined $\widetilde{Z}(\cdot,s)=\int_{0}^{\infty}~dt~e^{-st}~Z(\cdot,t)$ as the Laplace transform of the function $Z(\cdot,t)$. Eq. (2) can be used to obtain all the moments of the completion times. For instance, the MFPT can be defined as
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle\equiv\int_{0}^{\infty}dt~\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)=\frac{\langle T_{N}(L,x_{0})\rangle}{\epsilon_{0}(L,x_{0},N)},$ (3)
where $\langle T_{N}(L,x_{0})\rangle=\int_{0}^{\infty}~dt~q_{N}(L,x_{0},t)$ is the mean fastest first passage time out of $N$ searchers to reach either the target or the resetting threshold and $\epsilon_{0}(L,x_{0},N)=1-\int_{0}^{\infty}~dt~j_{L,N}(L,x_{0},t)$ is the splitting probability that any of the $N$ searchers reaches the target first before hitting the threshold.
Since the searchers are non-interacting and identical in nature, we can write $q_{N}(L,x_{0},t)=\left[Q(L,x_{0},t)\right]^{N}$, where $Q(L,x_{0},t)$ is the survival probability of a single searcher, starting at $x_{0}\in[0,L]$ upto time $t$, with absorbing boundaries at both 0 and $L$ Majumdar et al. (2020). On the other hand, $j_{L,N}(L,x_{0},t)$ consists of current to the resetting threshold from a single searcher i.e. $j_{L,1}(L,x_{0},t)$ barring the rest N1N-1 of them which survive with probability $\left[Q(L,x_{0},t)\right]^{N-1}$ so that $j_{L,N}(L,x_{0},t)=Nj_{L,1}(L,x_{0},t)\left[Q(L,x_{0},t)\right]^{N-1}$ (the factor $N$ accounts for the fact that any of the $N$ searchers can contribute to $j_{L,1}(L,x_{0},t)$). Substituting these statistical metrics into Eq. (3), we arrive at
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\frac{\int_{0}^{\infty}dt~[Q(L,x_{0},t)]^{N}}{1-\int_{0}^{\infty}dt~Nj_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}},$ (4)
which is a general expression for the MFPT and a key result of this paper. Importantly, the derivation does not rely on specific dynamics of the searchers or target configurations, and moreover, the framework can be generalized to higher-dimensional search processes.
Ballistic search under TR mechanism. To illustrate the theoretical framework developed above, we examine the paradigmatic case of ballistic searchers with randomized velocity. Such models have been commonly used to describe navigation pattern for some foragers such as microzooplankton, jackals and marine predators Reynolds and Bartumeus (2009); Sims et al. (2008), to scrutinize optimal strategies for identifying randomly located target sites Santos et al. (2004); James et al. (2008) and also in statistical physics Villarroel and Montero (2018); Biroli et al. (2024); Martens et al. (2012); Sprenger et al. (2023); Bechinger et al. (2016). Crucially, its analytical tractability offers valuable insights that enhance our understanding of TR-facilitated target search processes. Subsequently, we have extended our analysis to diffusive searchers.
We assume that each ballistic searcher moves with a random initial velocity drawn from a distribution $\phi(v)$. If the target is found a priori, the process therein ends. However, if the resetting threshold is encountered prior to the target, all the searchers are simultaneously reset to $x_{0}$ following which they draw new random velocities from the same distribution. As outlined above, we first derive the survival probability and the probability current for a single ballistic searcher (Sec. S1 in SI )
$Q(L,x_{0},t)$ $=\Phi\left(\frac{L-x_{0}}{t}\right)+\Phi\left(\frac{x_{0}}{t}\right),$ (5)
$~j_{L,1}(L,x_{0},t)$ $=\frac{L-x_{0}}{t^{2}}~\phi\left(\frac{L-x_{0}}{t}\right),$ (6)
where $\Phi(v)=\int_{0}^{v}dv^{\prime}\phi(v^{\prime})$; and substitute these into Eq. (4) to obtain the MFPT. For symmetric velocity distribution and $N=1$, this reads SI
$\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle=2L\int_{0}^{\infty}dv~\frac{\phi(v)}{v}.$ (7)
Remarkably, the MFPT does not depend on the starting point $x_{0}$ for any symmetric velocity distribution $\phi(v)$. Notably, the integral in Eq. (7) converges only when $\phi(v)\to 0$ as $v\to 0$ so that $\phi(v)/v$ becomes integrable around $v=0$. Simply put, if $\phi(v)$ does not vanish as $v\to 0$, there is a nonzero probability of observing realizations with searchers acquiring an almost vanishing velocity. These searchers will hit either the target or the resetting threshold in an exceedingly large timescale resulting in a divergent MFPT.
MFPT for multiple searchers ($N>1$): To continue the analysis for multiple searchers, we will assume that the velocities are chosen from a probability distribution given by $\phi(v)=\frac{1}{2v_{0}}e^{-|v|/v_{0}}$. While the MFPT for a single searcher is diverging, as evident from Eq. (7), the MFPT for the collective walkers e.g., $N\geq 2$ turns out to be a finite quantity. For the exponential kernel, the MFPT can be scaled as $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u\equiv\frac{x_{0}}{L})\rangle=\frac{v_{0}}{x_{0}}\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle={\cal F}\left(u,N\right)$, where $\mathcal{F}$ is the scaling function and depends only on $N$ and a dimensionless ratio $0\leq u\leq 1$. Computation of ${\cal F}(u,N)$ for $N>1$ is a challenging task that we undertake in SI , yielding
$\mathcal{F}(u,N)=\frac{\int_{0}^{\infty}dt\,\left[1-\frac{1}{2}\,e^{-u/t}-\frac{1}{2}\,e^{-(1-u)/t}\right]^{N}}{N\,u^{2}\,\int_{0}^{\infty}\frac{dt}{2t^{2}}\,e^{-u/t}\,\left[1-\frac{1}{2}\,e^{-u/t}-\frac{1}{2}\,e^{-(1-u)/t}\right]^{N-1}}\,.$ (8)
For $N=2$, a closed form expression can be obtained following a non-trivial computation (Sec. S3 in SI )
${\cal F}(u,2)=\frac{2}{u\,(3-2u)}\,\left[\ln 2-u\,\ln u-(1-u)\,\ln(1-u)\right],$ (9)
while for $N>2$, we can extract the following asymptotic forms which will be useful in later analysis
${\cal F}(u,N)\approx\begin{cases}&\hskip-8.5359pt\frac{a_{N}}{2^{N}-1}\,\frac{1}{u}\quad\quad\quad\quad\quad\quad\quad\quad\quad{\rm as}\quad u\to 0\,\vskip 5.69046pt\\ &\hskip-7.11317pta_{N}-N\,(1-u)\,\ln(1-u)\,\quad{\rm as}\quad u\to 1,\end{cases}$ (10)
where $a_{N}=\int_{0}^{\infty}\frac{dx}{x^{2}}\left(1-e^{-x}\right)^{N}$ is a constant. The scaling function ${\cal F}(u,N)$, evaluated as a function of $u$ for fixed $N$, diverges as $u\to 0$, and decreases monotonically with increasing $u$ when $N<4$. Interestingly, for $N\geq 4$, the function exhibits non-monotonic behavior: it first develops a local minimum, followed by a local maximum, and eventually approaches a global minimum as $u\to 1$, see Fig. 2(a) and a representative case for $N=7$ in Fig. 2(b) (solid curve).
Delving further, we observed that, in the absence of TR ($L\to\infty$ or equivalently $u=x_{0}/L\to 0$ limit), the FPT density decays as $\sim\frac{Nx_{0}}{2^{N}t^{2}}$, leading to a divergent MFPT for any $N$. In stark contrast, under optimal TR conditions at $u=1$, the distribution exhibits a faster decay, $f_{\mathcal{T}_{N}^{\text{TR}}}(x_{0},t)\sim t^{-N-1}$, ensuring a finite MFPT as $N\geq 2$ (see Sec. S8 for the derivation of the exact and universal FPT density that holds for arbitrary symmetric velocity distribution).
Optimization and limiting behavior of the MFPT.– To study the optimal behavior, we rewrite the MFPT by summing over the stochastic timescales
$\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle=\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle\times\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle+\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle,$~~ (11)
where $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle=\frac{1-\epsilon_{0}(u,N)}{\epsilon_{0}(u,N)}$ is the mean number of TR events upto the completion, and $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ is the mean conditional waiting time between consecutive TR events and $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ is the mean time for the first-passage to the target conditioned on no TR events. Numerical plots of these quantities are given in Fig. 2(b) (details in SI ).
In the limit $u\to 1$ when the resetting threshold is at the closest proximity to $x_{0}$, the resetting events are frequent and we find $\langle{\mathcal{N}}_{\text{TR}}(u\to 1,N)\rangle=2^{N}-1$ SI . However, since all such resetting events consume almost no time with $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u\to 1)\rangle\to 0$, the cumulative contribution to the MFPT from TR events becomes almost negligible. Notably, the finite contribution to the MFPT in this case comes solely from those correlated events where all the ballistic searchers collectively sample negative velocities right after the resetting or at the inception to drift right toward the target, marking a successful completion. Thus, $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u\to 1)\rangle$ effectively contributes to the MFPT whose asymptotic form is given by Eq. (10). The typical trajectories in this limit (see the bottom-right panel in Fig. 2b) confirm this observation. Quite interestingly, this emergent global optimality at $u\to 1$ aligns with a widely used threshold control strategy in chaos theory Sinha (2001); Bhowmick et al. (2014).
However, when the searchers start from a finite distance away from the resetting threshold ($u\lessapprox 1$), $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ become finite (recall that it is exactly zero for the $u\to 1$ limit) and $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$ remains large enough so that the first term in Eq. (11) dominates the MFPT leading to the local maxima observed in Fig. 2(b) (corresponding trajectories shown in the top-right figure for $u=0.9$). Further tuning initial configuration of the searchers in the direction of the target (i.e. reducing $u$ further) ensures $\epsilon_{0}>1/2$ (see Fig. (S2) in SI ) and subsequently, we see a gradual decrease in the MFPT upto a local minimum. Stochastic realizations at the bottom-left of Fig. 2(b) depict this scenario.
Moreover, in the limit $u\to 0$ where the threshold is far away from the starting position, each searcher can acquire a positive or negative velocity with probability $1/2$ for the symmetric velocity distribution and thus, there is a finite probability $\frac{1}{2^{N}}$ that all the $N$-searchers possess a positive velocity that is away from the target. The probability of selecting a unfavorable positive velocity, albeit vanishingly small, remains finite and yield in MFPT with a diverging form $\sim 1/u$ given by Eq. (10) (Sec. S5 in SI ). Contributing realizations (for instance, shown in the top-left panel in Fig. 2b for $u=0.0015$) corroborate this observation.
Finally, examining the MFPT with respect to $N$, we find that depending on $u$ and the interaction mediated by $N$, this variation can be either monotonic or non-monotonic. Nevertheless, for all $N$, the MFPT attains its minimum at $u=1$ (see Fig. S2), further supporting our claim in the ballistic search scenario.
Figure 3: Behavior of the cost function with respect to $u=x_{0}/L$ for $N=3$ and $\beta=1$. The cost attains a minimum at an optimal $u^{*}$ – the variation of which as a function of $N$ is shown in the inset panel. The squares represent data from simulations.
Cost of TR mechanism. A central quest in the resetting discipline is to harness the cost, be dynamic De Bruyne et al. (2020); Sunil et al. (2023, 2024) or thermodynamic Tal-Friedman et al. (2020, 2025); Olsen et al. (2024); Gupta et al. (2020); Pal et al. (2023), of the overall process. Imitating the penalty during a power grid breakdown Ajjarapu et al. (1994); Feng et al. (2000), we posit, for each resetting event, a cost that is proportional to the MFPT as well as the total number of TR events. Henceforth, the goal is to determine the optimal operation of this system, for which the cost function
$C_{N}(u)=\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle+\beta N\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle,$ (12)
is minimal. Here, $\beta$ is a constant that can be thought of cost per resetting event for a single particle and thus, we multiply by $N$ to enumerate the total cost for the “reset operation”. For the ballistic searchers, Fig. 3 shows that $C_{N}(u)$ attains a minimum for an optimal $u=u^{*}$ for a fixed $N$. Note that unlike the MFPT, the cost function is not globally minimized at $u\to 1$ since in this limit the resetting frequency is significantly high and the second term in Eq. (12) dominates. On the other hand, if the resetting threshold is far away from the starting position, the MFPT gradually increases, contributing to diverging cost. In the inset of Fig. 3, we show the variation of $u^{*}$ with respect to $N$, further noting that large number of searchers can systematically reduce the optimal distance $u^{*}\to 0$. Interestingly, the overall cost function remains optimized despite the increase in individual cost through $\beta$, accompanied by a gradual rise in $u^{*}$ SI . Furthermore, $u^{*}$ decreases with increasing $\beta$ for a fixed $N$, thus reflecting the high cost of frequent resets SI . Evidently, this collective cost optimization is different than that of the MFPT, indicating that a judicious choice of optimal $u$ can leverage to a decision based biasing between the MFPT and the cost function.
Conclusions. Motivated by threshold-driven phenomena that refer to processes or events that occur only when a system surpasses a critical threshold, we study the first passage properties of an arbitrary search dynamics under threshold resetting mechanism. We develop a general theoretical framework and study the optimization properties of the MFPT which is a useful metric in learning the efficiency of a target search process. The optimization reported here also resonates with the performance maximization by the agents (robots or foraging animals) that collaborate and achieve tasks (for instance finding the desired target between many competitive ones by a fish school in Couzin et al. (2011)) in an optimal way.
In a classical foraging model where the searchers move ballistically, we discover a non-trivial optimization that suggests that persistent resetting to the threshold can render the MFPT globally minimum – an effect reminiscent to the threshold mechanism in nonlinear chaotic systems Sinha (2001); Bhowmick et al. (2014). Extending the analysis to $N$ diffusive searchers also reveals rich optimization behaviors of the MFPT SI ; Biswas et al. (2025). By defining a cost function for such systems, we further demonstrate that the collective TR mechanism can achieve optimal cost.
To extend the formalism to capture more realistic scenarios, we can incorporate an overhead time for information transfer. This can be done in two ways: First, we can add a random (or deterministic) information passing time $\mathcal{T}^{O}$ whenever a searcher hits the threshold, after which all of them are reset to the initial position. In this case, one can show that the overall MFPT of the system would increase (however, this computation excludes the time required to transfer information to the others once a searcher successfully locates the target) Biswas et al. (2025). A more involved extension would be to consider space–time coupled return protocols, where after a TR event searchers return to the initial configuration via finite-time spatial dynamics Pal et al. (2020); Tal-Friedman et al. (2020); Biswas et al. (2024).
Finally, it is worth noting that the formalism can also be applied to extract higher order moments of the search time in higher dimensions and to the dynamics of non-coordinate variables like a fluctuating potential as in the case of firing neurons where the membrane potential of an individual synapse can reset after reaching a critical value Burkitt (2006); Taillefumier and Magnasco (2013). Ongoing works also include characterizing nonequilibrium stationary states for such long-range systems under TR Biswas et al. (2025); Biroli et al. (2025). Given its simple implementation yet achieving target functionality makes the TR mechanism a promising avenue in the target search problem which can have far reaching implications beyond statistical physics to foraging, control theory, finance, structural failure-repair analysis and in computing systems.
Acknowledgments. SNM thanks M. Biroli, S. Redner and G. Schehr for discussions on related models. AP sincerely thanks D. Ghosh for pointing towards some relevant works. The numerical calculations reported in this work were carried out on the Kamet cluster, which is maintained and supported by the Institute of Mathematical Science’s High-Performance Computing Center. AB and AP gratefully acknowledge research support from the Department of Atomic Energy, Government of India. SNM acknowledges support from ANR Grant No. ANR- 23- CE30-0020-01 EDIPS. SNM and AP thank the Higgs Center for Theoretical Physics, Edinburgh, for hospitality during the workshop “New Vistas in Stochastic Resetting” and Korea Institute for advanced Study(KIAS), Seoul, for hospitality during the conference “Nonequilibrium Statistical Physics of Complex Systems” where several discussions related to the project took place. SNM and AP also acknowledge the research support from the International Research Project (IRP) titled “Classical and quantum dynamics in out of equilibrium systems” by CNRS, France.
Supplemental material for
“Target search optimization by threshold resetting”
This Supplemental Material provides detailed derivations of some of our main results highlighted in the Letter. Moreover, it also provides additional discussions that support our finding as announced in the Letter.
S1 Survival probability and probability current of a single ballistic searcher
In this section, we compute the survival probability $Q(L,x_{0},t)$ and the current through the resetting threshold $j_{L,1}(L,x_{0},t)$ for a single ballistic searcher as used in Eq. (5) and (6) of the main text. If $v_{i}$ denotes the random initial velocity of the $i^{th}$ searcher then it can survive up to time $t$ without hitting either target or resetting threshold in two ways: First, if $v_{i}t<L-x_{0}$ when $v_{i}>0$ so that the walker starts moving towards the resetting threshold but failing to reach the same in the given time and second, when $|v_{i}|t<x_{0}$ if the initial velocity $v_{i}<0$ so that the walker starts moving towards the target yet fails to hit it by time $t$. Summing over these two possibilities, we can write the survival probability as
$Q(L,x_{0},t)=\int_{0}^{\infty}dv~\phi(v)~\theta(L-x_{0}-vt)+\int_{-\infty}^{0}dv~\phi(v)~\theta(x_{0}-|v|t),$ (S1)
where $\theta(z)$ is the Heaviside theta function which takes values unity only when $z>0$, and 0 otherwise. Assuming $\phi(v)$ to be symmetric so that $\phi(-v)=\phi(v)$, one can rewrite the above equation as
$Q(L,x_{0},t)=\int_{0}^{\frac{L-x_{0}}{t}}dv~\phi(v)+\int_{0}^{\frac{x_{0}}{t}}dv~\phi(v).$ (S2)
Denoting the function $\Phi(v)$ as
$\Phi(v)=\int_{0}^{v}dv^{\prime}\phi(v^{\prime}),$ (S3)
we finally arrive at
$Q(L,x_{0},t)=\Phi\left(\frac{L-x_{0}}{t}\right)+\Phi\left(\frac{x_{0}}{t}\right),$ (S4)
which is Eq. (5) of the main text. The probability current through the resetting threshold at $x=L$ is essentially the total number of accumulated particles there at time $t$ so that
$j_{L,1}(L,x_{0},t)$ $=\int_{0}^{\infty}dv~\phi(v)~\delta\left(t-\frac{L-x_{0}}{v}\right)$
$=\frac{L-x_{0}}{t^{2}}~\phi\left(\frac{L-x_{0}}{t}\right),$ (S5)
which is Eq. (6) of the main text. Here, $\delta(x)$ is the Dirac-delta function. Note that only the positive velocities contribute to the current through the threshold while the negative velocities contribute to the current through the target at the origin given by
$j_{0,1}(L,x_{0},t)$ $=\int_{-\infty}^{0}dv~\phi(v)~\delta\left(t-\frac{x_{0}}{v}\right)$
$=\frac{x_{0}}{t^{2}}~\phi\left(\frac{x_{0}}{t}\right).$ (S6)
S2 MFPT of a single ballistic searcher under TR Mechanism Eq. (7)
This section extends the derivation of the MFPT of a single searcher under TR mechanism which was announced in Eq. (7) of the main text. To this end, we start by deriving the numerator in Eq. (3) of the main text by setting $N=1$ which yields the unconditional MFPT for a single particle to reach either the resetting threshold or the target
$\langle T_{1}(L,x_{0})\rangle$ $=\int_{0}^{\infty}dt~Q(L,x_{0},t)$
$=\int_{0}^{\infty}dt\int_{0}^{\infty}dv~\phi(v)~\theta(L-x_{0}-vt)+\int_{0}^{\infty}dt\int_{-\infty}^{0}dv~\phi(v)~\theta(x_{0}-|v|t)$
$=\int_{0}^{\infty}dv~\phi(v)\int_{0}^{\infty}dt~\theta(L-x_{0}-vt)+\int_{-\infty}^{0}dv~\phi(v)~\int_{0}^{\infty}dt\theta(x_{0}-|v|t)$
$=\int_{0}^{\infty}dv~\phi(v)~\left(\frac{L-x_{0}}{v}\right)+\int_{-\infty}^{0}dv~\phi(v)~\left(\frac{x_{0}}{|v|}\right).$ (S7)
Using the symmetry property of the velocity distribution i.e. $\phi(-v)=\phi(v)$, we can further simplify the above equation to find
$\langle T_{1}(L,x_{0})\rangle$ $=\int_{0}^{\infty}dv~\phi(v)~\left(\frac{L-x_{0}}{v}\right)+\int_{0}^{\infty}dv~\phi(v)~\left(\frac{x_{0}}{v}\right)$
$=L\int_{0}^{\infty}dv~\frac{\phi(v)}{v}.$ (S8)
Let us turn our attention to compute the denominator of Eq. (3) from the main text for $N=1$. We first note that it is just the splitting probability $\epsilon_{0}(L,x_{0},1)$ of a single searcher to reach the target located at the origin before hitting the threshold. For a symmetric velocity distribution, $\epsilon_{0}(L,x_{0},1)$ is simply $1/2$ since any searcher can acquire a positive (negative) velocity with probability $1/2$ indifferent to the initial coordinate $x_{0}$. This can also be seen formally as
$\epsilon_{0}(L,x_{0},1)$ $=1-\int_{0}^{\infty}dt~j_{L,1}(L,x_{0},t)$
$=\int_{0}^{\infty}dt~j_{0,1}(L,x_{0},t)$
$=\int_{0}^{\infty}dt~\frac{x_{0}}{t^{2}}~\phi(x_{0}/t)$
$=\frac{1}{2}.$ (S9)
Putting together both the numerator and denominator in Eq. (3) of the main text, the MFPT under TR mechanism for a single searcher takes the form
$\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle=\frac{\langle T_{1}(L,x_{0})\rangle}{\epsilon_{0}(L,x_{0},1)}=2L\int_{0}^{\infty}dv~\frac{\phi(v)}{v},$ (S10)
which is Eq. (7) of the main text.
S3 Derivation of the scaling function ${\cal F}(u,N)$ as in Eq. (8)
In this section, we derive an exact form of the scaling function ${\cal F}(u,N)$ which was announced in the main text. To this end, we use Eq. (4) from the main text and note that we need the single particle survival probability (as in Eq. (S4)) and the probability current to the resetting threshold (as in Eq. (S5)) for the exponential velocity distribution $\phi(v)=\frac{1}{2v_{0}}e^{-|v|/v_{0}}$. In this case, we have $\Phi(v)=\frac{1}{2}\left(1-e^{-v/v_{0}}\right)$. Defining the following variables
$u=x_{0}/L,~~\tau_{b}=L/v_{0},$ (S11)
where $\tau_{b}$ is the timescale of a ballistic walker with speed $v_{0}$ and using Eq. (S4) and Eq. (S5), we can now compute the survival probability and the currents to the resetting threshold and the target respectively which take the following form
$Q(u,t)=1-\frac{1}{2}\exp\left({-\frac{\tau_{b}(1-u)}{t}}\right)-\frac{1}{2}\exp\left({-\frac{\tau_{b}u}{t}}\right),$ (S12)
$j_{L,1}(u,t)=\frac{\tau_{b}(1-u)}{2t^{2}}\exp\left({-\frac{\tau_{b}(1-u)}{t}}\right),$ (S13)
$j_{0,1}(u,t)=\frac{\tau_{b}u}{2t^{2}}\exp\left({-\frac{\tau_{b}u}{t}}\right).$ (S14)
We now recall from the main text that the splitting probability for any of the $N$ particles to hit the target is given by
$\epsilon_{0}(L,x_{0},N)$ $=1-\int_{0}^{\infty}dt~j_{L,N}(L,x_{0},t)$
$=1-\int_{0}^{\infty}dt~Nj_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}$
$=\int_{0}^{\infty}dt~Nj_{0,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}.$ (S15)
Plugging the results from Eq. (S12)-(S15) into Eq. (4) of the main text, we find
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\frac{\int_{0}^{\infty}dt~\left[1-\frac{1}{2}\exp\left({-\frac{[1-u]\tau_{b}}{t}}\right)-\frac{1}{2}\exp\left({-\frac{\tau_{b}u}{t}}\right)\right]^{N}}{\int_{0}^{\infty}dt~N\left[\frac{\tau_{b}u}{2t^{2}}\exp\left({-\frac{u\tau_{b}}{t}}\right)\right]\left[1-\frac{1}{2}\exp\left({-\frac{[1-u]\tau_{b}}{t}}\right)-\frac{1}{2}\exp\left({-\frac{\tau_{b}u}{t}}\right)\right]^{N-1}}.$ (S16)
Now we make the change of variable as $t\to\frac{\tau_{b}}{t}$ to obtain
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\frac{\tau_{b}\int_{0}^{\infty}dt\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N}}{Nu\int_{0}^{\infty}dt\left[\frac{1}{2t^{2}}\exp\left({-\frac{u}{t}}\right)\right]\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}}.$ (S17)
Noting $\tau_{b}=\frac{1}{u}\frac{x_{0}}{v}$ from Eq. (S11) and substituting into the above expression for MFPT, we find
$\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\frac{\frac{x_{0}}{v}\int_{0}^{\infty}dt\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N}}{Nu^{2}\int_{0}^{\infty}dt\left[\frac{1}{2t^{2}}\exp\left({-\frac{u}{t}}\right)\right]\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}},$
$\implies$ $\mathcal{F}(u,N)=\frac{v_{0}}{x_{0}}\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\frac{\int_{0}^{\infty}dt\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N}}{Nu^{2}\int_{0}^{\infty}dt\left[\frac{1}{2t^{2}}\exp\left({-\frac{u}{t}}\right)\right]\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}},$ (S18)
which is essentially Eq. (8) of the main text. In Fig. (S1), we show the variation of the scaling function with $u$ for different $N$. The scaling function ${\cal F}(u,N)$, evaluated as a function of $u$ for fixed $N$, diverges as $u\to 0$, and decreases monotonically with increasing $u$ when $N<4$. Interestingly, for $N\geq 4$, the function exhibits non-monotonic behavior: it first develops a local minimum, followed by a local maximum, and eventually approaches a global minimum as $u=1$.
Figure S1: Variation of the scaling function $\mathcal{F}(u,N)$ as in Eq. (S18) with respect to $u$ for different choices of $N$. The markers represents the results from numerical simulation which match perfectly with the theoretical curves indicated by the solid lines.
In Fig. (S2) we also show the variation of the MFPT with respect to $N$ for several choices of $u$. From the figure it is evident that (i) the MFPT for any $u<1$ can be monotonic or non-monotonic in nature with respect to $N$, however (ii) the envelope for MFPT corresponding to $u=1$ remains the lowest for any choice of $N$. The results reconcile with the key result for the ballistic searchers that $u=1$ is the optimal value for the collective ballistic search. It also implies that the long-range interaction reinforced by the collective resetting of $N$ searchers works in the most optimal way when $u=1$.
Figure S2: Variation of the MFPT of ballistic searchers under TR with respect to $N$ for various values of $u$. Note that for $u=1$ the MFPT is lowest for any values of $N$ as shown by the red line. The markers represents results from numerical simulation.
S4 Derivation of the closed form for the scaling function $\mathcal{F}(u,N)$ for $N=2$ as in Eq. (9)
Here we derive the scaling function (or the scaled MFPT) for $N=2$ ballistic searchers i.e. $\mathcal{F}(u,2)$. It is convenient to make the change of variable $t=1/x$ and rewrite the scaling function (Eq. (8) of the main text or Eq. (S18)) as
${\cal F}(u,N)=\frac{2}{N\,u^{2}}\,\frac{I_{N}(u)}{J_{N}(u)}\,$, (S19)
where $I_{N}(u)$ and $J_{N}(u)$ are given by
$I_{N}(u)$ $=$ $\int_{0}^{\infty}\frac{dx}{x^{2}}\,\left[1-\frac{1}{2}\,e^{-u\,x}-\frac{1}{2}\,e^{-(1-u)x}\right]^{N}$ (S20)
$J_{N}(u)$ $=$ $\int_{0}^{\infty}dx\,e^{-u\,x}\,\left[1-\frac{1}{2}\,e^{-u\,x}-\frac{1}{2}\,e^{-(1-u)x}\right]^{N-1}\,.$ (S21)
For general $N\geq 2$, it is not easy to compute the two integrals $I_{N}(u)$ and $J_{N}(u)$, and hence ${\cal F}(u,N)$ in Eq. (S19). However for $N=2$, as we show below, one can compute ${\cal F}(u,2)$ exactly for all $u\in[0,1]$. This will serve also as a benchmark to compare to the asymptotic behaviors of ${\cal F}(u,N)$ for general $N$ in the two asymptotic limits $u\to 0$ and $u\to 1$.
For $N=2$, let us start with the integral $J_{2}(u)$ in Eq. (S21) which is relatively simpler to evaluate. Setting $N=2$ in Eq. (S21) we get
$J_{2}(u)=\int_{0}^{\infty}dx\,e^{-ux}\,\left[1-\frac{1}{2}\,e^{-u\,x}-\frac{1}{2}\,e^{-(1-u)x}\right]=\frac{3}{4u}-\frac{1}{2}\,.$ (S22)
The computation of $I_{2}(u)$ in Eq. (S20) turns out to be tricky as we show in below. Following Eq. (S20), we write
$I_{2}(u)=\int_{0}^{\infty}\frac{dx}{x^{2}}\left[1-\frac{1}{2}\,e^{-u\,x}-\frac{1}{2}\,e^{-(1-u)\,x}\right]^{2}\,.$ (S23)
To get rid of the singularity $x^{-2}$ in the integrand, it is convenient to first take a derivative with respect to $u$. This gives
$I_{2}^{\prime}(u)=\frac{dI_{2}(u)}{du}=\int_{0}^{\infty}\frac{dx}{x}\left[1-\frac{1}{2}\,e^{-u\,x}-\frac{1}{2}\,e^{-(1-u)\,x}\right]\,\left[e^{-u\,x}-e^{-(1-u)\,x}\right]\,.$ (S24)
We note one important fact from this equation that will be useful later, namely
$I_{2}^{\prime}\left(u=\frac{1}{2}\right)=0\,.$ (S25)
The integrand in Eq. (S24) still has an singularity $1/x$. To get rid of this term, let us take one more derivative with respect to $u$. Further simplification leads to the following integral that can be trivially computed to yield
$I_{2}^{\prime\prime}(u)=\frac{d^{2}I_{2}(u)}{du^{2}}=\int_{0}^{\infty}dx\,\left[e^{-2ux}-e^{-2(1-u)x}-e^{-ux}-e^{-(1-u)x}\right]=-\frac{1}{2u}-\frac{1}{2(1-u)}\,.$ (S26)
Now, integrating it back with respect to $u$ gives
$I_{2}^{\prime}(u)=-\frac{1}{2}\ln u+\frac{1}{2}\ln(1-u)+C_{1}\,$, (S27)
where the unknown constant $C_{1}=0$ is fixed from the condition in Eq. (S25). Thus we get
$I_{2}^{\prime}(u)=-\frac{1}{2}\ln u+\frac{1}{2}\ln(1-u)\,.$ (S28)
Figure S3: Comparison of the scaling function for $N=2$ as in Eq. (S39) with that obtained via numerical integration of Eq. (S18) .
Integrating once more we get
$I_{2}(u)=-\frac{1}{2}\left[u\ln u-u\right]+\frac{1}{2}\left[1-u-(1-u)\ln(1-u)\right]+C_{0}\,$, (S29)
where the constant $C_{0}$ is fixed from the value $I_{2}(0)$ in Eq. (S23). Thus we finally get
$I_{2}(u)=I_{2}(0)-\frac{1}{2}\,u\,\ln u-\frac{1}{2}\,(1-u)\,\ln(1-u)\,$, (S30)
where, using Eq. (S23),
$I_{2}(0)=\frac{1}{4}\,a_{2}\,\quad{\rm with}\quad a_{2}=\int_{0}^{\infty}\frac{dx}{x^{2}}\left[1-e^{-ux}\right]^{2}\,.$ (S31)
To evaluate the integral $a_{2}$, we do one integration by part that gives
$a_{2}=2\int_{0}^{\infty}\frac{dx}{x}\,e^{-x}\left(1-e^{-x}\right)\,.$ (S32)
To do the remaining integral in Eq. (S32), let us introduce an auxiliary parameter $b$ and define
$a_{2}(b)=2\,\int_{0}^{\infty}\frac{dx}{x}\,e^{-x}\left(1-e^{-b\,x}\right)\,$, (S33)
such that $a_{2}(b=1)=a_{2}$. Differentiating Eq. (S33) with respect to $b$ we get
$a_{2}^{\prime}(b)=\frac{da_{2}(b)}{db}=2\,\int_{0}^{\infty}dx\,e^{-(b+1)\,x}=\frac{2}{b+1}\,.$ (S34)
Integrating it back with respect to $b$ and using $a_{2}(b=0)=0$, we get
$a_{2}(b)=2\,\ln(b+1)\,.$ (S35)
Consequently,
$a_{2}=a_{2}(b=1)=2\,\ln 2\,.$ (S36)
Using this result in Eq. (S31) gives
$I_{2}(0)=\frac{a_{2}}{4}=\frac{1}{2}\,\ln 2\,.$ (S37)
Following Eq. (S30), we get the complete expression for $I_{2}(u)$ as
$I_{2}(u)=\frac{1}{2}\,\left[\ln 2-u\,\ln u-(1-u)\,\ln(1-u)\right]\,.$ (S38)
Substituting $J_{2}(u)$ from Eq. (S22) and $I_{2}(u)$ from Eq. (S38) into Eq. (S19) with $N=2$ gives the exact scaling function ${\cal F}(u,2)$ as
${\cal F}(u,2)=\frac{2}{u\,(3-2u)}\,\left[\ln 2-u\,\ln u-(1-u)\,\ln(1-u)\right]\,$, (S39)
as mentioned in Eq. (9) of the main text. In Fig. (S3) we show that the exact result as in Eq. (S39) matches exactly with the numerical integration of Eq. (S18) for $N=2$. The limiting behaviors of ${\cal F}(u,2)$ are given by
${\cal F}(u,2)\approx\begin{cases}\frac{2\ln 2}{3\,u}\,\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad{\rm as}\quad u\to 0\\ \\ 2\,\ln 2-2\ (1-u)\,\ln(1-u)\,\,\,\quad{\rm as}\quad u\to 1\,.\end{cases}$ (S40)
Derivation of the scaling function for general $N$ is subject matter of the next section.
S5 Asymptotic behaviors of ${\cal F}(u,N)$ for general $N\geq 2$ as in Eq. (10)
We now consider the scaling function ${\cal F}(u,N)$ in Eq. (S19) for general $N\geq 2$. Unlike the $N=2$ case, it turns out to be much harder to compute the scaling function explicitly for general $N>2$. However, as mentioned earlier, one can extract the asymptotic behaviors of ${\cal F}(u,N)$, for arbitrary $N\geq 2$, in the two limits $u\to 0$ and $u\to 1$ as shown below.
S5.1 The limit $u\to 0$
We will show that for general $N\geq 2$, the scaling function ${\cal F}(u,N)$ diverges as $u\to 0$, and this leading divergence originates from the denominator $J_{N}(u)$ in Eq. (S19) as $u\to 0$. Hence, for the numerator $I_{N}(u)$ in Eq. (S19), it is enough only the leading $u=0$ behavior where it turns out to be finite for all $N\geq 1$. This becomes evident by setting $u=0$ in Eq. (S20), and it gives
$I_{N}(0)=\frac{a_{N}}{2^{N}}\,\quad{\rm where}\quad a_{N}=\int_{0}^{\infty}\frac{dx}{x^{2}}\left(1-e^{-x}\right)^{N}\,.$ (S41)
We then need to evaluate $J_{N}(u)$ in Eq. (S21) in the limit $u\to 0$. By naively putting $u=0$ in Eq. (S21) one finds that the resulting integral diverges. To extract the leading divergence of $J_{N}(u)$ as $u\to 0$, we first make a change of variable ux=yu\,x=y in the integral in Eq. (S21). This gives
$J_{N}(u)=\frac{1}{u}\,\int_{0}^{\infty}dy\,e^{-y}\,\left[1-\frac{1}{2}\,e^{-y}-\frac{1}{2}\,e^{-\frac{(1-u)}{u}\,y}\right]^{N-1}\,.$ (S42)
Now, in the limit $u\to 0$, we can safely neglect the third term inside the parenthesis $\ldots$] since the resulting integral is convergent. This gives the leading order behavior as $u\to 0$
$J_{N}(u)\approx\frac{1}{u}\,\int_{0}^{\infty}dy\,e^{-y}\,\left[1-\frac{1}{2}\,e^{-y}\right]^{N-1}\,.$ (S43)
Making the change of variable $z=1-\frac{1}{2}\,e^{-y}$, the integral can be done exactly leading to
$J_{N}(u)\approx\left[\frac{2}{N}\left(1-2^{-N}\right)\right]\,\frac{1}{u}\,\quad{\rm as}\quad u\to 0\,.$ (S44)
Substituting Eq. (S44) and Eq. (S41) into the exact expression for ${\cal F}(u,N)$ in Eq. (S19) then gives the leading small $u$ behavior
${\cal F}(u,N)\approx\frac{a_{N}}{2^{N}-1}\,\frac{1}{u}\quad{\rm as}\quad u\to 0\,$, (S45)
where $a_{N}$ is defined in Eq. (S41). Note that this result is valid for any $N\geq 2$. For $N=2$, using $a_{2}=2\,\ln 2$ from Eq. (S36), we see that the general result in Eq. (S45) for $N=2$ matches perfectly with the leading asymptotic behavior as $u\to 0$ derived in Eq. (S40) from the exact scaling function for $N=2$.
S5.2 The limit $u\to 1$
We now consider the $u\to 1$ limit of the scaling function ${\cal F}(u,N)$ in Eq. (S19). It is easier to first consider the denominator $J_{N}(u)$ defined in Eq. (S21). Setting $u=1$, we find
$J_{N}(1)=\frac{1}{2^{N-1}}\,\int_{0}^{\infty}dx\,e^{-x}\,\left(1-e^{-x}\right)^{N-1}=\frac{1}{N\,2^{N-1}}\,.$ (S46)
The last integral is trivially performed upon the change of variable $z=1-e^{-x}$. In fact, setting $u=1-\epsilon$, it is easy to compute the next order term for small $\epsilon$. One finds that
$J_{N}(1-\epsilon)=\frac{1}{N\,2^{N-1}}+b_{N}\,\epsilon+O\left(\epsilon^{2}\right)\,,\quad{\rm where}\quad b_{N}=-\frac{1}{2^{N-1}}\,\int_{0}^{1}dz\,z^{N-1}\,\ln(1-z)\,.$ (S47)
We will see later that the $O(\epsilon)$ term in $J_{N}(1-\epsilon)$ does not contribute to the first two leading behaviors of ${\cal F}(u,N)$ as $u\to 1$. We now turn to the integral $I_{N}(u)$ defined in Eq (S20). Setting $u=1-\epsilon$ we get
$I_{N}(1-\epsilon)=W_{N}(\epsilon)=\int_{0}^{\infty}\frac{dx}{x^{2}}\,\left[1-\frac{1}{2}e^{-(1-\epsilon)\,x}-\frac{1}{2}\,e^{-\epsilon\,x}\right]^{N}\,.$ (S48)
Now, exactly at $u=1$, i.e., at $\epsilon=0$, the integral is convergent and we get
$I_{N}(1)=W_{N}(0)=\frac{a_{N}}{2^{N}}\,,\quad{\rm where}\quad a_{N}=\int_{0}^{\infty}\frac{dx}{x^{2}}\left(1-e^{-x}\right)^{N}\,.$ (S49)
Note that the same $a_{N}$ also appeared in Eq. (S41) near the limit $u\to 0$ which is not surprising given the fact that the function $I_{N}(u)$ defined in Eq. (S20) is symmetric around $u=1/2$.
Our goal is to evaluate the next order correction to $I_{N}(1-\epsilon)=W_{N}(\epsilon)$ as $\epsilon\to 0$. To do this, we first differentiate $W_{N}(\epsilon)$ in Eq. (S49) once with respect to $\epsilon$. This gives
$W_{N}^{\prime}(\epsilon)=\frac{N}{2}\int_{0}^{\infty}\frac{dx}{x}\,\left[1-\frac{1}{2}\,e^{-(1-\epsilon)\,x}-\frac{1}{2}\,e^{-\epsilon\,x}\right]^{N-1}\,\left[e^{-\epsilon\,x}-e^{-(1-\epsilon)\,x}\right]\,.$ (S50)
We note that as $\epsilon\to 0$, the first derivative $W_{N}^{\prime}(\epsilon)$ also diverges. To extract this divergence, we take one more derivative of $W_{N}^{\prime}(\epsilon)$ in Eq. (S50) with respect to $\epsilon$. This gives
$W_{N}^{\prime\prime}(\epsilon)$ $=$ $\frac{N(N-1)}{4}\,\int_{0}^{\infty}dx\,\left[1-\frac{1}{2}\,e^{-(1-\epsilon)\,x}-\frac{1}{2}\,e^{-\epsilon\,x}\right]^{N-2}\,\left[e^{-\epsilon\,x}-e^{-(1-\epsilon)\,x}\right]^{2}$ (S51)
- $\frac{N}{2}\,\int_{0}^{\infty}dx\,\left[1-\frac{1}{2}\,e^{-(1-\epsilon)\,x}-\frac{1}{2}\,e^{-\epsilon\,x}\right]^{N-1}\,\left[e^{-\epsilon\,x}+e^{-(1-\epsilon)\,x}\right]\,.$ (S52)
Next we make a change of variable $\epsilon\,x=y$. This gives
$W_{N}^{\prime\prime}(\epsilon)$ $=$ $\frac{N(N-1)}{4\,\epsilon}\,\int_{0}^{\infty}dy\,\left[1-\frac{1}{2}\,e^{-y}-\frac{1}{2}\,e^{-\frac{(1-\epsilon)}{\epsilon}\,y}\right]^{N-2}\,\left[e^{-y}-e^{-\frac{(1-\epsilon)}{\epsilon}\,y}\right]^{2}$ (S53)
- $\frac{N}{2\,\epsilon}\,\int_{0}^{\infty}dy\,\left[1-\frac{1}{2}\,e^{-y}-\frac{1}{2}\,e^{-\frac{(1-\epsilon)}{\epsilon}\,y}\right]^{N-1}\,\left[e^{-y}+e^{-\frac{(1-\epsilon)}{\epsilon}\,y}\right]\,.$ (S54)
At this stage, we can now take the $\epsilon\to 0$ limit inside the integral and neglect the term $e^{-\frac{1-\epsilon}{\epsilon}\,y}$ everywhere, since the resulting integrals remain convergent. This leads to the leading $\epsilon\to 0$ behavior of the second derivative
$W_{N}^{\prime\prime}(\epsilon)\approx\frac{N(N-1)}{4\,\epsilon}\,\int_{0}^{\infty}dy\,e^{-2\,y}\,\left[1-\frac{1}{2}\,e^{-y}\right]^{N-2}-\frac{N}{2\,\epsilon}\,\int_{0}^{\infty}dy\,e^{-y}\,\left[1-\frac{1}{2}\,e^{-y}\right]^{N-1}\,.$ (S55)
These two integrals can now be easily performed by making the change of variable $1-\frac{1}{2}\,e^{-y}=z$. This gives a very simple expression for the leading behavior as $\epsilon\to 0$
$W_{N}^{\prime\prime}(\epsilon)\approx-\frac{N}{2^{N}\,\epsilon}\,.$ (S56)
Integrating back once with respect to $\epsilon$ gives
$W_{N}^{\prime}(\epsilon)\approx-\frac{N}{2^{N}}\,\ln(\epsilon)+C_{3}\,$, (S57)
where $C_{3}$ is an integration constant whose value is not relevant for us. Integrating further with respect to $\epsilon$ we get the two leading order terms as $\epsilon\to 0$
$I_{N}(1-\epsilon)=W_{N}(\epsilon)=W_{N}(0)-\frac{N}{2^{N}}\,\epsilon\,\ln(\epsilon)+O(\epsilon)\,.$ (S58)
Using $W_{N}(0)$ from Eq. (S49), we then obtain, as $u\to 1$,
$I_{N}(u)=\frac{a_{N}}{2^{N}}\,-\frac{N}{2^{N}}\,(1-u)\,\ln(1-u)+O(1-u)\,$, (S59)
where $a_{N}$ is given in Eq. (S49).
Finally, substituting the leading behaviors of $I_{N}(u)$ from Eq. (S59) and of $J_{N}(u)$ from Eq. (S47) as $u\to 1$ into our main formula (S19), we obtain the following asymptotic behavior of ${\cal F}(u,N)$ as $u\to 1$ for arbitrary $N\geq 2$
${\cal F}(u,N)=a_{N}-N\,(1-u)\,\ln(1-u)+O(1-u)\,.$ (S60)
Hence, summarizing, the two limiting asymptotic behaviors of the scaling function ${\cal F}(u,N)$, for arbitrary $N\geq 2$, are given by
[…]
where we recall that the constant $a_{N}$ is defined in Eq. (S49). Note that for $N=2$, using $a_{2}=2\,\ln 2$ from Eq. (S36), the general result in Eq. (S61) does coincide with the leading asymptotic behaviors of ${\cal F}(u,2)$ in Eq. (S40) obtained from the exact scaling function for $N=2$.
S6 Derivations of $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle,\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ and $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$
The MFPT under TR mechanism as in Eq. (4) of the main text can be interpreted as a sum over stochastic timescales emanating from all possible events eventually leading to a completion of the process. Following Eq. (11) of the main text, we recall
$\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle=\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle\times\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle+\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle,$ (S62)
where individual quantities were defined in the main text. In this section, we explicitly derive these quantities and elaborate further.
S6.1 Derivation of $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$
A stochastic trajectory under TR mechanism can experience multiple resetting events before the first passage to the target. The number of these TR events fluctuates between the realizations and our aim is to compute the mean number of resetting events till the first passage. Since, the system can undergo only two possible events: TR events (which can be called as a unsuccessful event) and a first passage event (a successful event), clearly the number of TR events is distributed according to a geometric distribution. Here, the success probability is nothing but the splitting probability to the threshold $\epsilon_{L}(u,N)$ so that the unsuccess probability is given by $\epsilon_{0}(u,N)$ which is the splitting probability to the target. Thus, the mean number of TR events can be obtained as the mean of the geometric distribution
$\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$ $=\sum_{k=0}^{\infty}k(\epsilon_{L}(u,N))^{k}\times\epsilon_{0}(u,N))=\frac{1-\epsilon_{0}(u,N)}{\epsilon_{0}(u,N)},$ (S63)
where $\epsilon_{0}(u,N)$ can be found using Eq. (S15) as follows
$\epsilon_{0}(u,N)=1-\epsilon_{L}(u,N)=Nu\int_{0}^{\infty}d\tau~\frac{e^{-\frac{u}{\tau}}}{2\tau^{2}}\left(1-\frac{1}{2}e^{-\frac{1-u}{\tau}}-\frac{e^{-\frac{u}{\tau}}}{2}\right)^{N-1}.$ (S64)
In Fig. (S4) we show the variation of the splitting probabilities with $u$ for $N=7$. Note that for $u<1/2$ we have $\epsilon_{0}>1/2$. In Fig. (S5)(a), we plot $\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle$ as a function of $u$ for several values of $N$.
Figure S4: Variation of the splitting probabilities with $u$ for $N=7$ as found from Eq. (S64). Note that when $u<1/2$ we have $\epsilon_{0}>\epsilon_{L}$.
S6.2 Derivation of $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$
Recall that the random variable ${\mathbb{T}}_{N}^{\text{TR}}$ denotes the waiting time between two consecutive TR events conditioned on survival. The normalized density of this conditional time can then be written as
$f_{{{\mathbb{T}}_{N}^{\text{TR}}}}(t)$ $=\frac{j_{L,N}(L,x_{0},t)}{\epsilon_{L}(L,x_{0},N)}$
$=\frac{j_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}}{\int_{0}^{\infty}dt~j_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}},$ (S65)
where the common factor $N$ has been canceled out from the numerator and denominator. From the distribution, one can find the mean time between two consecutive resetting events as
$\langle{\mathbb{T}}_{N}^{\text{TR}}(L,x_{0})\rangle=\int_{0}^{\infty}dt~tf_{{{\mathbb{T}}_{N}^{\text{TR}}}}(t).$ (S66)
Plugging the values of the survival probability as in Eq. (S12) and the probability current to the threshold as in Eq. (S13) in the above equation one arrives at $\langle{{\mathbb{T}}_{N}^{\text{TR}}}\rangle$ (in a dimensionless form)
$\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ $=\frac{v_{0}}{x_{0}}\langle{{\mathbb{T}}_{N}^{\text{TR}}}(L,x_{0})\rangle=\frac{\int_{0}^{\infty}dt~\frac{1}{2t}\exp\left({-\frac{1-u}{t}}\right)\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}}{u\int_{0}^{\infty}dt~\frac{1}{2t^{2}}\exp\left({-\frac{1-u}{t}}\right)\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}}.$ (S67)
In Fig. Fig. (S5)(b), we plot $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ as a function of $u$ for different $N$.
Figure S5: Variation of $\langle{\mathcal{N}}_{\text{TR}}(u,N)$ as in Eq. (S63), $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u)\rangle$ as in Eq. (S67) and $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ as in Eq. (S69) with respect to $u$ for distinct values of $N$.
S6.3 Derivation of $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$
Finally, we turn our attention to the random variable ${\mathbb{T}}_{N}^{\text{FP}}$ that denotes the random time for the first passage since the last TR event. This is simply the conditional time that the target is found before the threshold by any of the $N$ searchers. Following these arguments, we can write the normalized density of ${\mathbb{T}}_{N}^{\text{FP}}$ as
$f_{{{\mathbb{T}}_{N}^{\text{FP}}}}(t)$ $=\frac{j_{0,N}(L,x_{0},t)}{\epsilon_{0}(L,x_{0},N)}$
$=\frac{j_{0,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}}{\int_{0}^{\infty}dt~j_{0,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}}.$ (S68)
Finally plugging the survival probability Eq. (S12) and the probability current to the target as in Eq. (S14) we find the mean of ${\mathbb{T}}_{N}^{\text{FP}}$ (in the dimensionless form) as
$\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ $=\frac{v_{0}}{x_{0}}\int_{0}^{\infty}dt~tf_{{\mathbb{T}}_{N}^{\text{FP}}}(t)=\frac{\int_{0}^{\infty}dt~\frac{1}{2t}\exp\left({-\frac{u}{t}}\right)\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}}{u\int_{0}^{\infty}dt~\frac{1}{2t^{2}}\exp\left({-\frac{u}{t}}\right)\left[1-\frac{1}{2}\exp\left({-\frac{1-u}{t}}\right)-\frac{1}{2}\exp\left({-\frac{u}{t}}\right)\right]^{N-1}}.$ (S69)
In Fig. (S5)(c), we plot $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ as a function of $u$ for different values of $N$.
S6.4 Asymptotic behavior in the limit $u\to 1$
In the following, we present the asymptotic forms of the above-mentioned observables that was used in the main text to explain the behavior of MFPT in the limit of $u\to 1$. For instance, the splitting probability to the target, in this limit, takes the following form
$\epsilon_{0}(u=1,N)=N\int_{0}^{\infty}d\tau~\frac{e^{-\frac{1}{\tau}}}{2\tau^{2}}\left(\frac{1}{2}-\frac{e^{-\frac{1}{\tau}}}{2}\right)^{N-1}=\frac{1}{2^{N}}.$ (S70)
Thus, from Eq. (S63), we have
$\langle{\mathcal{N}}_{\text{TR}}(u=1,N)\rangle=2^{N}-1.$ (S71)
The mean conditional time between TR events, however, in this limit vanishes to zero as can be verified by setting $u=1$ in Eq. (S67) as seen below
$\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u\to 1)\rangle$ $=\frac{\int_{0}^{\infty}dt~\frac{1}{t}\left[1-\exp\left({-\frac{1}{t}}\right)\right]^{N-1}}{\int_{0}^{\infty}dt~\frac{1}{t^{2}}\left[1-\exp\left({-\frac{1}{t}}\right)\right]^{N-1}}\to 0.$ (S72)
On the other hand, the conditional mean first passage time to the target $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u)\rangle$ has a finite value as can be seen from Eq. (S69) by setting the limit $u\to 1$ as follows
$\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u\to 1)\rangle=N\int_{0}^{\infty}dt~\frac{1}{t}\exp\left({-\frac{1}{t}}\right)\left[1-\exp\left({-\frac{1}{t}}\right)\right]^{N-1},$ (S73)
which was used to analyze the form of the MFPT in the $u\to 1$ limit.
S7 Diffusive searchers under TR
In this section, we extend the discussion on the first-passage properties of diffusive search under TR. The formalism prescribed in the main text can be directly used to compute the statistical properties for this process as well. We start by recalling the single searcher diffusive propagator $G(x,t|x_{0})$ in the presence of two absorbing boundaries at $x=0$ and $x=L$ from literature Redner (2001)
$G(x,t|x_{0})=\frac{2}{L}\sum_{n=0}^{\infty}\sin\left(\frac{\pi nx}{L}\right)\sin\left(\frac{\pi nx_{0}}{L}\right)e^{-\frac{n^{2}\pi^{2}Dt}{L^{2}}},$ (S74)
where $D$ is the diffusion constant. Next, we would need to compute the single searcher survival probability $Q(L,x_{0},t)$, the flux through the origin $j_{0,1}(L,x_{0},t)$ and the threshold $j_{L,1}(L,x_{0},t)$ respectively which can be obtained using the following relations
$Q(L,x_{0},t)=\int_{0}^{L}G(x,t|x_{0})dx,$ (S75)
$j_{0,1}(L,x_{0},t)=D\left.\frac{\partial G(x,t|x_{0})}{\partial x}\right|_{x=0},$ (S76)
$j_{L,1}(L,x_{0},t)=-D\left.\frac{\partial G(x,t|x_{0})}{\partial x}\right|_{x=L}.$ (S77)
Using the exact expression of the propagator for the diffusive searchers, we find
$Q(u,t)=\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (S78)
$j_{0,1}(u,t)=\frac{2\pi}{\tau_{d}}\sum_{n=1}^{\infty}n\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (S79)
$j_{L,1}(u,t)=-\frac{2\pi}{\tau_{d}}\sum_{n=1}^{\infty}n(-1)^{n}\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (S80)
where
$u=\frac{x_{0}}{L},~~~~~\tau_{d}=\frac{L^{2}}{D}.$ (S81)
Evidently, the dimensionless MFPT under TR should become a function of the parameters $u$ and $N$, so that
$\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle=\frac{D}{x_{0}^{2}}\langle\mathcal{T}_{N}^{\text{TR}}(L,x_{0})\rangle=\mathcal{F}(u,N),$ (S82)
with $\mathcal{F}(u,N)$ being the scaling function. Substituting the relevant observables for the single searcher as obtained in Eq. (S78)-(S80) into Eq. (4) of the main text, we obtain the exact expression of the scaling function $\mathcal{F}(u,N)$ given by
$\mathcal{F}(u,N)$ $=\frac{1}{u^{2}}\left(\frac{\int_{0}^{\infty}dt~\left[\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-n^{2}\pi^{2}t}\right]^{N}}{N\int_{0}^{\infty}dt~\left[2\pi\sum_{n=1}^{\infty}n\sin\left(n\pi u\right)e^{-n^{2}\pi^{2}t}\right]\left[\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-n^{2}\pi^{2}t}\right]^{N-1}}\right).$ (S83)
Through the above change of variables, we have recast the dimensionless mean first-passage time (MFPT), $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle$, as a function of a single dimensionless parameter $u\in(0,1)$, effectively eliminating dependence on the original parameters $L$, $x_{0}$, and $D$. While the expression in Eq. (S83) does not admit a closed-form solution for arbitrary $N$, it can be evaluated numerically with high precision using Mathematica, yielding results in excellent agreement with numerical simulations (Fig. S6a). In what follows, we examine in detail the behavior of the scaled MFPT—i.e., the scaling function $\mathcal{F}(u,N)$—as a function of both $u$ and $N$, and highlight the physical insights that emerge from this analysis.
Figure S6: (a) Variation of the non-dimensionalized MFPT $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle$ as a function of $u=x_{0}/L$. The solid lines represent the analytical result as in Eq. (S83) and the markers represent the simulation results for different values of number of searchers NN. The dashed horizontal lines represent the reset-free case i.e. when $L\to\infty$ or $u\to 0$. Note that for $N=1$ and $N=2$ the respective dashed lines are not shown since the MFPT of the underlying process in these cases are infinity. In the simulation, we have fixed $x_{0}=0.5$ and varied $L$. The value of the diffusion constant was set at $D=1$. (b) Shows the variation of the point $u_{\text{opt}}$ (inset shows an example for the case of $N=3$ searchers) with respect to $N$. Note that, there are two possible ways to realize both the figures Fig. S6(a,b) physically. Either one can fix $x_{0}$ and vary $L$ from $x_{0},\infty]$ or one can fix $L$ and vary $x_{0}$ from $(0,L)$. Both the cases, although, physically different, will indeed yield the same curves as shown in the above figures. This is because, the dimensionless MFPT in Eq. (S83) does not depend explicitly on $x_{0}$ or $L$, but only on the ratio $u=x_{0}/L$. From Fig. S6(b) we see that $u_{\text{opt}}\to 0$ as $N\to\infty$.
S7.1 Optimization with respect to threshold
As before, we ask the following question: Can the variation of $u$ lead to a minimal MFPT? To explore this, we plot the MFPT as a function of $u$ in Fig. S6(a) for various values of $N$. Remarkably, for all $N\geq 1$, the MFPT exhibits a clear minimum at an optimal value of $u$. Let us discuss these cases in the following.
Single diffusive searcher $(N=1)$: For a single diffusive searcher, from Eq. (S83) the MFPT is found to be
$\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle=\frac{\frac{L^{2}}{2D}(x_{0}/L)(1-x_{0}/L)}{1-x_{0}/L}=\frac{x_{0}L}{2D}$
or, $\langle\overline{\mathcal{T}}_{1}^{\text{TR}}(u)\rangle=\frac{D}{x_{0}^{2}}\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle=\frac{1}{2u}.$ (S84)
Here, the numerator of $\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle$ is the unconditional MFPT for a single Brownian searcher to reach either the boundary at $x=0$ or $x=L$ Redner (2001). The denominator is the splitting probability for it to reach the target Redner (2001). From Eq. (S84) it is evident that the MFPT for a single diffusive searcher decreases as $u^{-1}$ as also shown in Fig. S6(a). It is the lowest when $L\to x_{0}$ (i.e., $u\to 1$) where $\langle\overline{\mathcal{T}}_{1}^{\text{TR}}(u=1)\rangle=1/2$. Clearly here $u\to 1$ is the optimal point as also observed in Fig. S6(b). Physically, such behavior can be explained in the following way: The threshold effectively biases the searcher’s motion towards the target by resetting it from $x_{0}$ whenever it goes away from the target. Keeping $x_{0}$ fixed, as one decreases $L$, therefore increasing $u$, the chances of the searchers wandering away from the target also diminish, resulting in a lower MFPT.
Multiple diffusive searchers $(N\geq 2)$: The MFPT exhibits quite distinct features for $N\geq 2$ than that for $N=1$. In here, the MFPT curves show a non-monotonic behaviour with respect to $u$ as seen in Fig. S6(a).
The limit $u\to 0$ is just the reset-free case where the threshold is kept at infinity (assuming $x_{0}$ to be fixed). In this case, the MFPT $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u\to 0)\rangle$ is just the mean fastest first passage time out of $N$ searchers to reach the target in the absence of the threshold. After some simplifications, the MFPT in this limit can be found to be
$\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u\to 0)\rangle=\frac{1}{2}\int_{0}^{\infty}dy~y~\left[\text{erf}\left(\frac{1}{y}\right)\right]^{N},$ (S85)
which is a finite quantity only for $N\geq 3$. The value of the MFPT in the limit ($u\to 0$) as in Eq. (S85) is shown by dashed horizontal lines in Fig. S6(a) for $N\geq 3$ (recall that for $N\leq 2$, the underlying search has infinite MFPT and thus TR naturally expedites). As one increases $u$, the MFPT starts to decrease, showing a minimum at some intermediate value $u=u_{\text{opt}}$ so that $\left.\frac{\partial\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle}{\partial u}\right|_{u=u_{\text{opt}}}=0$ (explicitly shown in the inset of Fig. S6(b) for $N=3$) before increasing again as $u$ approaches unity. As stressed earlier, $u_{\text{opt}}$ can be thought of as the analog of optimal resetting rate in the traditional resetting set-up. In Fig. S6(b) we show the variation of $u_{\text{opt}}$ with respect to $N$. For $N=1$ we find $u_{\text{opt}}=1$ since the lowest MFPT occurs only at $u=1$ as evident from Eq. (S84). However, for any other values of $N\geq 2$ we find $0<u_{\text{opt}}<1$.
For intermediate values of $u$, the MFPT curves fall below their respective dashed lines—each representing the reset-free case for different NN—clearly indicating that threshold resetting enhances search efficiency [Fig. S6(a)]. Although it is known that the MFPT for diffusive searchers is finite only for $N\geq 3$ Lindenberg et al. (1980); Krapivsky et al. (2010), the TR mechanism enables a further reduction by optimally constraining the searchers’ trajectories. This efficiency gain arises from the presence of a finite threshold that prevents searchers from diffusing beyond the latter. However, as $u\to 1$, the MFPT begins to increase again for all $N\geq 2$, in contrast to the monotonic behavior observed for $N=1$. In this regime, the initial position $x_{0}$ is very close to the threshold $L$, resulting in a high splitting probability $\epsilon_{L}$ of reaching the threshold rather than the target. Consequently, frequent resets disrupt the search process. As $N$ increases, the likelihood that at least one searcher reaches the threshold before the target also increases, leading to a suppression of successful target-hitting events and a corresponding rise in the MFPT.
S8 First-passage time density under TR
In this section, we present a detailed study of the FPT density of the collective search in the absence and presence of threshold resetting. The large time behavior of the FPT density becomes particularly useful to understand and compare the first-passage time moments between the processes.
S8.1 FPT distribution of the ballistic searchers without TR ($u\to 0$)
The survival probability for a single ballistic searcher (starting from $x_{0}$ and target at $x=0$) in absence of the threshold is given by
$Q_{0}(x_{0},t)$ $=\int_{0}^{\infty}dv~\phi(v)+\int_{-\infty}^{0}dv~\phi(v)~\theta(x_{0}-|v|t)$
$=\frac{1}{2}+\int_{0}^{\frac{x_{0}}{t}}dv~\phi(v)$
$=\frac{1}{2}+\Phi\left(\frac{x_{0}}{t}\right).$ (S86)
Thus, the survival probability for $N$ independent ballistic searchers are given by
$Q_{0}^{N}(x_{0},t)=\left[Q_{0}(x_{0},t)\right]^{N}=\left[\frac{1}{2}+\Phi\left(\frac{x_{0}}{t}\right)\right]^{N},$ (S87)
which can be used to deduct the FPT distribution as follows
$f_{T_{0}}(x_{0},t)$ $=-\frac{\partial Q_{0}^{N}(x_{0},t)}{\partial t}$
$=N\left[\frac{1}{2}+\Phi\left(\frac{x_{0}}{t}\right)\right]^{N-1}\left[\frac{x_{0}}{t^{2}}\phi\left(\frac{x_{0}}{t}\right)\right].$ (S88)
For the exponential velocity distribution i.e., $\phi(v)=\frac{1}{2v_{0}}e^{-v/v_{0}}$ we have $\Phi(v)=\frac{1}{2}(1-e^{-v/v_{0}})$. Finally plugging this expressions into the above results (setting $v_{0}=1$ without loss of generality) we have
$f_{T_{0}}(x_{0},t)=N\left(1-\frac{1}{2}e^{-x_{0}/t}\right)^{N-1}\left(\frac{x_{0}}{2t^{2}}e^{-x_{0}/t}\right).$ (S89)
In the limit $t\to\infty$ it has the following asymptotic behavior
$f_{T_{0}}(x_{0},t\to\infty)\sim\frac{Nx_{0}}{2^{N}t^{2}}.$ (S90)
The exponent of $t$ at asymptotic limit is 2 (i.e., $N$ independent) leading to a diverging MFPT for any given values of $N$. The variation of the FPT density with $t$ is shown in Fig. S7(a) which shows an excellent agreement with numerical simulations.
Figure S7: (a) The scaled FPT density $f_{T_{0}}(x_{0},t)/\frac{Nx_{0}}{2^{N}}$ for the ballistic searchers (without TR) exhibits $\sim 1/t^{2}$ decay for any values of $N$. The markers represents results from numerical simulation. The solid black line is the asymptotic result as in Eq. (S90). (b) The FPT density $f_{\mathcal{T}_{N}^{\text{TR}}}(L=x_{0},t)$ under optimal TR. Note that the tail exponent becomes NN-dependent in this case as mentioned in the main text – see Eq. (S98).
S8.2 FPT distribution of the ballistic searchers under optimal TR
We have noted from the main text that the optimal MFPT occurs for the ballistic case when $u=1$ so that $L=x_{0}$. In here, we provide an exact derivation for the FPT distribution of the ballistic searchers at this optimality. To this end, let us recall the renewal equation for the survival probability as in Eq. (1) of the main text
$\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t)=q_{N}(L,x_{0},t)+\int_{0}^{t}dt^{\prime}j_{L,N}(L,x_{0},t^{\prime})\mathcal{Q}_{N}^{\text{TR}}(L,x_{0},t-t^{\prime}).$~~~~ (S91)
At the optimal case of $u=1$ or $L=x_{0}$, the particles can survive in two possible ways. First is the the no-resetting case in which the survival probability is simply $q_{N}(L=x_{0})=[Q(L,t)]^{N}$ (also see the main text). In the other possible scenario, one or multiple resetting events can happen. This implies that at least one of the searchers will attain a positive velocity so that resetting will surely take place – the probability of such occurrence is given by $1-\frac{1}{2^{N}}$. Since the resetting here takes no time to occur as $u=1$, the probability current is simply given $j_{L,N}(L=x_{0},t^{\prime})=\left(1-\frac{1}{2^{N}}\right)\delta(t^{\prime})$, where $\delta(t^{\prime})$ is the Dirac-delta function. Combining both these possibilities, the renewal equation for the survival probability at $L=x_{0}$ takes the form
$\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t)=[Q(L,t)]^{N}+\left(1-\frac{1}{2^{N}}\right)\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t).$ (S92)
Rearranging, we obtain an exact expression for the survival probability at the optimality i.e. $L=x_{0}$ as
$\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t)=[2Q(L,t)]^{N},$ (S93)
which is a very interesting result since it holds for arbitrary symmetric velocity distribution of the ballistic searchers. From Eq. (5) of the main text we find that the single particle survival probability $Q(L=x_{0},t)$ for a symmetric velocity distribution is given by
$Q(L,t)=\Phi\left(\frac{L}{t}\right),$ (S94)
where $\Phi(v)=\int_{0}^{v}dv\phi(v)$. Thus the survival probability has the final form
$\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t)=\left[2\Phi\left(\frac{L}{t}\right)\right]^{N}.$ (S95)
Recall that, for the exponential velocity distribution i.e., $\phi(v)=\frac{1}{2v_{0}}e^{-v/v_{0}}$ we have $\Phi(v)=\frac{1}{2}(1-e^{-v/v_{0}})$. Thus the survival probability in this case has the form (setting $v_{0}=1$ without any loss of generality)
$\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t)=\left(1-e^{-L/t}\right)^{N},$ (S96)
from which the FPT distribution can be obtained via the relation $f_{\mathcal{T}_{N}^{\text{TR}}}(L=x_{0},t)=-\frac{\partial\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t)}{\partial t}$ and is given by
$f_{\mathcal{T}_{N}^{\text{TR}}}(L=x_{0},t)=\frac{NL}{t^{2}}e^{-L/t}\left(1-e^{-L/t}\right)^{N-1},$ (S97)
with the following asymptotic behavior
$f_{\mathcal{T}_{N}^{\text{TR}}}(L=x_{0},t)\sim NL^{N}t^{-N-1}~~\text{as}~~t\to\infty.$ (S98)
Thus the distribution of the FPT distribution under optimal TR for the ballistic walkers with exponential velocities follows a power tail $\sim t^{-N-1}$. This exponent makes sure that the MFPT is finite immediately when $N\geq 2$. This result is verified against numerical simulations and is shown in Fig. S7(b).
Alternative derivation: One can also derive the FPT distribution as in Eq. (S97) but starting from Eq. (2) of the main text. This derivation, however, relies upon the exponential velocity distribution of the ballistic searchers. Let us recall Eq. (2) in the following
$\widetilde{\mathcal{Q}}_{N}^{\text{TR}}(L,x_{0},s)=\frac{\int_{0}^{\infty}dt~e^{-st}[Q(L,x_{0},t)]^{N}}{1-\int_{0}^{\infty}dt~e^{-st}Nj_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}}.$ (S99)
For the ballistic walkers with exponential velocity distribution we earlier derived
$Q(L,x_{0},t)=1-\frac{1}{2}\left[e^{-x_{0}/t}+e^{-(L-x_{0})/t}\right]$ (S100)
$j_{L,1}(L,x_{0},t)=\frac{L-x_{0}}{2t^{2}}e^{-(L-x_{0})/t}.$ (S101)
In terms of these quantities the denominator can be written as
$\mathcal{D}\equiv 1-N\int_{0}^{\infty}dt~e^{-st}N\frac{L-x_{0}}{2t^{2}}e^{-(L-x_{0})/t}\left[1-\frac{1}{2}(e^{-x_{0}/t}+e^{-(L-x_{0})/t})\right]^{N-1}.$ (S102)
To extract the limit $L\to x_{0}$ we first rescale $t$ as $t=(L-x_{0})y$. In terms of the rescaled variable $y$ the integral can be written as
$\mathcal{D}=1-\frac{N}{2}\int_{0}^{\infty}\frac{dy}{y^{2}}e^{-1/y}e^{-s(L-x_{0})y}\left[1-\frac{1}{2}e^{-\frac{x_{0}}{(L-x_{0})y}}-\frac{1}{2}e^{-1/y}\right]^{N-1}.$ (S103)
Now we can take the limit $L\to x_{0}$ to have
$\mathcal{D}=1-\frac{N}{2}\int_{0}^{\infty}\frac{dy}{y^{2}}e^{-1/y}e^{-s(L-x_{0})y}\left[1-\frac{1}{2}e^{-1/y}\right]^{N-1},$ (S104)
that can be exactly solved to give the result
$\mathcal{D}=\frac{1}{2^{N}},$ (S105)
which is also the splitting probability $\epsilon_{0}(u=1,N)$ derived earlier [see Eq. (S70) in the SM]. Substituting $\mathcal{D}$ into Eq. (S99), we have
$\widetilde{\mathcal{Q}}_{N}^{\text{TR}}(L=x_{0},s)=2^{N}\int_{0}^{\infty}dt~e^{-st}[Q(L=x_{0},t)]^{N}=\int_{0}^{\infty}dt~e^{-st}[1-e^{-L/t}]^{N},$ (S106)
which is already in the form of a Laplace transform. Thus, by inspection, the survival probability in time domain can be written as
$\mathcal{Q}_{N}^{\text{TR}}(L=x_{0},t)=\left(1-e^{-L/t}\right)^{N},$ (S107)
which is Eq. (S96) as derived earlier.
S8.3 FPT distribution of the diffusive searchers in the absence ($u\to 0$) and presence of TR
For diffusive searchers, the FPT distribution is known to have the following asymptotic behavior in the absence of TR Lindenberg et al. (1980); Krapivsky et al. (2010)
$f_{T_{0}}(x_{0},t\to\infty)\sim t^{-1-\frac{N}{2}}.$ (S108)
Note that the exponent is $N$ dependent in this case and renders an finite MFPT when $N>2$. However, as the TR is introduced, the FPT densities for various $N$ at their respective optimal $u_{\text{opt}}$ are shown to exhibit exponential tails $\sim e^{-at}$, ensuring finite moments–refer to Fig. S8.
Figure S8: Numerical FPT density for diffusion with TR for various NN demonstrating an exponential decay.
S9 Variation of cost function with respect to $\beta$
We recall the definition of cost-function from the main text
$C_{N}(u)=\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle+\beta N\langle{\mathcal{N}}_{\text{TR}}(u,N)\rangle,$ (S109)
where $\beta$ is a constant that can be thought of cost per resetting event for a single searcher and thus, we multiply by $N$ to enumerate the total cost for the “reset operation”. To understand the variation of the cost function with respect to $\beta$, we plot $C_{N}(u)$ for different $u$ and find that the curves exhibit non-monotonic dependence on $u$ for any $\beta$ – see Fig. S9(a). It is further noted that $u^{*}$ monotonically decreases as $\beta$ is varied to larger values – see Fig. S9(b). Overall, this study further strengthens our claim that threshold resetting can indeed optimize the collective cost regardless of the individual increase in cost. In Fig. S10 we show the variation of $u^{*}$ with $N$, but for several $\beta$. We particularly observe that, for a given $N$, optimal $u^{*}$ decreases as $\beta$ is increased. Physically, for a higher value of $\beta$ in this regime, the cost per resetting event becomes large so that resetting the system too frequently (i.e., increasing $u$, see Fig. S5(a)) will significantly increase the cost function. Thus, the optimal way for the system to function is to have a lower value of $u$ to avoid high cost due to threshold resetting.
Figure S9: Panel (a): Variation of the cost function with respect to $u$ for different $\beta$ for the ballistic search case. The curves maintain non-monotonicity with a finite optimal $u^{*}$. Panel (b): Dependence of optimal $u^{*}$ as a function of $\beta$.
Figure S10: Variation of $u^{*}$ (where the cost function is optimal) as a function of $N$ for different values of $\beta$.
Figure S11: A higher dimension embedding depiction to realize simultaneous resetting for two searchers in 1D by mapping to a single searcher in 2D. A single 2D searcher starts from $(x_{0},x_{0})$ inside a square with the target shown by the dashed lines and the resetting threshold by the solid lines. The searcher moves with certain velocity $\vec{v}$, but as soon as it hits one of the thresholds, it is reset back to $(x_{0},x_{0})$. In this way, both the coordinates are reset simultaneously and new $\vec{v^{\prime}}$ is drawn. This is equivalent to 2 independent searchers getting reset at once when one of them hits the threshold and subsequently, they restart with new velocities. Here, we have shown two subsequent resetting events followed by a successful target search with a renewed velocity $\vec{v^{\prime\prime}}$.